### v1

# Grammatical Error Correction for Low-Resource Languages: The Case of Zarma

Abstract
Grammatical error correction (GEC) is important for improving written materials for low-resource languages like Zarma—spoken by over 5 million people in West Africa. Yet it remains a challenging problem. This study compares rule-based methods, machine translation (MT) models, and large language models (LLMs) for GEC in Zarma. We evaluate each approach’s effectiveness on our manually-built dataset of over 250,000 examples using synthetic and human-annotated data. Our experiments show that the MT-based approach using the M2M100 model outperforms others, achieving a detection rate of 95.82% and a suggestion accuracy of 78.90% in automatic evaluations, and scoring 3.0 out of 5.0 in logical/grammar error correction during MEs by native speakers. The rule-based method achieved perfect detection (100%) and high suggestion accuracy (96.27%) for spelling corrections but struggled with context-level errors. LLMs like MT5-small showed moderate performance with a detection rate of 90.62% and a suggestion accuracy of 57.15%. Our work highlights the potential of MT models to enhance GEC in low-resource languages, paving the way for more inclusive NLP tools.
Grammatical Error Correction for Low-Resource Languages: The Case of Zarma
Mamadou K. KEITA1, Christopher Homan3, Sofianou Abdoulaye Hamani1, Adwoa Bremang2, Marcos Zampieri4, Habibatou Abdoulaye Alfari2, Elysabhete Amadou Ibrahim2, Dennis Owusu2 1NILab, 2Ashesi University, 3Rochester Institute of Technology, 4George Mason University
1 Introduction
Grammatical error correction (GEC) is an essential task in natural language processing (NLP) that aims to improve the quality and readability of texts by correcting grammatical errors. GEC are important for enhancing written materials, significantly impacting educational outcomes, professional opportunities, and access to information. This is particularly relevant in under-resourced settings where limited access to academic resources and formal training can amplify language proficiency disparities. Furthermore, communities in these settings often rely on local languages for communication and knowledge transmission. While GEC tools are well-developed for high-resource languages—especially English—where extensive annotated datasets and standardized writing systems are available (Napoles et al., 2019), GEC presents significant challenges for low-resource languages.
Zarma, a Nilo-Saharan language spoken by over 5 million people across Niger and neighboring countries (Lewis et al., 2016), exemplifies the difficulties faced by many low-resource languages. The lack of standardized orthography and the limited availability of annotated data make it challenging to develop practical GEC tools. These challenges are not unique to Zarma but are shared by a broad class of low-resource languages, including many African and indigenous languages worldwide. Addressing these challenges requires innovative approaches that can work with minimal data and accommodate non-standardized texts.
The emergence of large language models (LLMs) gave notice to a new era in NLP, empowering machine learning models to learn and generate human-like text across many languages (Devlin et al., 2018; Brown et al., 2020). LLMs exhibit remarkable zero-shot and few-shot learning capabilities (Wan et al., 2023), which are potentially beneficial for low-resource languages. These capabilities allow LLMs to perform effectively with minimal data, making them valuable for languages with limited annotated datasets. However, the use of LLMs is also challenging in low-resource settings. They are primarily trained on data from high-resource languages, which may limit their performance when applied to low-resource languages. The lack of representative training data can lead to errors and biases, reducing their effectiveness. Additionally, the significant computational resources required to fine-tune and deploy LLMs can be a barrier in resource-constrained environments.
This study investigates the potential of LLMs and traditional models to improve GEC for Zarma by comparing conventional rule-based methods, machine translation (MT)-based models, and the novel application of LLMs for Zarma GEC. To explore the performance of these models on other low-resource languages, we replicate our Zarma GEC experiments with Bambara, a West African language spoken in Mali. Through a case study, we highlight the strengths and weaknesses of each method, aiming to bridge the linguistic gap in NLP for low-resource languages like Zarma and Bambara.
Our research addresses the following questions:
RQ1:
Do state-of-the-art LLM models outperform conventional rule-based and MT-based models on GEC for Zarma texts?
RQ2:
What are the specific strengths and limitations of each GEC approach in low-resource settings, considering the lack of standardized orthography and limited annotated data?
The main contributions of this paper are as follows:
  • A comprehensive evaluation of three distinct GEC approaches for Zarma (rule-based, MT-based, and LLM-based). Our findings show that the MT-based approach delivered the highest accuracy with a detection rate of 95.82%, a suggestion accuracy of 78.90%, and acceptable performance in zero-shot scenarios.
  • Reproduction of the experiments with Bambara—a manding family language—to confirm replicability and broaden the study’s scope beyond Zarma.
  • Development and planned public release (upon acceptance) of models fine-tuned for GEC of the tested languages.
2 Related Work
Cissé and Sadat (2023) present an approach for spellchecking Wolof, a language primarily spoken in Senegal. Their algorithm combines a dictionary lookup with an edit distance metric Levenshtein distance algorithm (Levenshtein, 1966) to identify and correct spelling errors.
Vydrin and collaborators developed Daba, a software package for grammar and spellchecking in Manding languages (Vydrin, 2014). Their work employs a rule-based system focusing on morphological analysis, addressing the agglutinative nature of Manding languages.
Researchers have explored the use of LLMs for language-specific tasks, including GEC, demonstrating their adaptability beyond high-resource languages. For instance, a study by Palma Gomez et al. (2023) showed that the MT5 model (Xue et al., 2020), a multilingual transformer model pre-trained on a massive dataset of text and code, could be effectively fine-tuned for GEC in Ukrainian. Song et al. (2023) present an innovative application of LLMs—specifically GPT-4 (OpenAI, 2023)—for generating explanations for grammatical errors. Another promising approach to GEC leverages pre-trained multilingual MT models. The study by (Luhtaru et al., 2024) introduced the "No Error Left Behind" approach, which uses models like MT5 and NLLB (Team et al., 2022). The fine-tuning process involves adapting the pre-trained MT model to treat error correction as a "translation" task, where the source language is the incorrect sentence and the target language is the corrected sentence. However, the research also identified a fundamental challenge: the trade-off between precision and recall when training with synthetic data. This finding shows the need for further research to optimize these models for low-resource GEC tasks, potentially through improved data augmentation techniques or developing specialized architectures for error correction.
3 Proposed Methods
3.1 GEC with LLMs
LLMs have significantly advanced NLP, exhibiting capabilities in multitasking, few-shot and multilingual learning. These models, extensively pre-trained on diverse datasets, demonstrate a remarkable ability to grasp nuanced aspects of language, reasoning, and context (Brown et al., 2020; Raffel et al., 2020). This section outlines our methodology for leveraging LLMs to develop a GEC tool for Zarma. The proposed GEC tool is designed to function independently of predefined grammar rules or lexicons, using the models’ few-shot learning capabilities for enhanced efficiency in low-resource scenarios.
3.1.1 Implementation
We employed two distinct approaches for LLM-based GEC:
Instruction and Error Explanation Fine-tuning
This method involves embedding training data within a contextual sentence structure, enhancing the model’s reasoning abilities, and facilitating effective learning from the examples. We use the following instruction prompt:
Prompt: "Zarma sentence: [Incorrect Sentence],Correct the zarma sentence: [Correct Sentence]Error Causes: : [Error $\begin{array}[]{l}\text{{Prompt}: "Zarma sentence: [{Incorrect Sentence}],}\\ \text{Correct the zarma sentence: [{Correct Sentence}]}\\ \text{{Error Causes: }: [{Error Cause}]."}\end{array}$
This format, particularly the error explanation component, is important for aiding the model’s comprehension of Zarma’s grammatical contexts and patterns, as demonstrated in previous research (Schick and Schütze, 2021; Wei et al., 2021).
Non-Prompt Fine-tuning Using Aligned Sentences
This approach involves fine-tuning the model directly on parallel data of incorrect and correct Zarma sentences without explicit prompts. This leverages the model’s ability to learn the implicit mapping between incorrect and correct forms.
3.2 GEC with MT Models
Exploring MT models for GEC represents a promising avenue for addressing GEC challenges in low-resource languages like Zarma. MT models, particularly those pre-trained on multilingual datasets, offer strong capabilities for processing text across diverse linguistic frameworks.
3.2.1 Implementation
We chose the M2M100 model (Fan et al., 2020) for its demonstrated ability to translate between many languages, indicating its potential to capture cross-linguistic patterns relevant to GEC. To adapt M2M100 for Zarma GEC, we fine-tuned it on a corrupted corpus designed to reflect common errors in Zarma text. This corpus was generated by applying a custom noise script, described in Section 4. The script introduces various errors, ensuring the training data effectively represents realistic challenges in real-world Zarma text. This corrupted corpus, paired with the original correct sentences, serves as the training data for M2M100, enabling the model to learn the mapping from incorrect to correct Zarma. Detailed training settings and evaluation metrics for this MT-based GEC approach can be found in Tables 7, 4, and 5.
3.3 Rule-based GEC for Zarma
As a baseline, we designed a rule-based GEC process, based on Levenshtein distance and the Bloom filter (Bloom, 1970), for Zarma (Figure 1). Additionally, we implemented a tool and API in Python. The tool—the first of its kind for Zarma, to our knowledge—is designed to cater to a wide range of users. It offers command-line and graphical user interfaces (GUI) and is much less computationally intensive than the LLM-based approach. Moreover, our results show that it provides a competitive spell correction performance compared to the LLMs- and MT-based approaches. To ensure the tool’s accessibility, we plan a public release upon acceptance.
Figure 1: Rule-Based GEC tool Workflow
The GEC process for Zarma text follows a structured sequence of steps, starting with the input text $T$, passing through a Bloom filter decision $B$ to determine if the word exists in the dictionary, leading to either a grammatical check or a correction algorithm, followed by a final dictionary lookup and correction $C$, as needed:
$C(T)=\begin{cases}T&\text{if }D(B(T))=\text{True},\\ \text{Correct}(T)&\text{if }D(B(T))=\text{False}.\end{cases}$ (1)
The detailed process includes:
Text Processing
The system begins by processing the text, where the Zarma content is segmented into words, punctuation, and spaces using regular expressions. This initial step ensures that every part of the text is ready for individual examination—a necessity given Zarma’s linguistic intricacies.
Bloom Filter
The Bloom filter—known for its space and time efficiency—performs probabilistic checks on whether a word might be in the dictionary. For a given word $w$, it uses a set of hash functions to probe various positions in a bit array. If all checked positions are flagged, $w$ might be in the dictionary:
$B(w)=\bigwedge_{i=1}^{k}\text{bit}[h_{i}(w)],$ (2)
If $B(T)=\text{False}$, the system assumes the word is not in the dictionary and immediately moves to the Correction Algorithm to suggest a correction. This process ensures efficiency when dealing with unknown words.
Dictionary Lookup
At the core of the system is the trie-based dictionary lookup. If the Bloom Filter decision is positive ($B(T)=\text{True}$), the system proceeds to a Grammatical Check to verify if the word is grammatically correct. If the word is found to be grammatically incorrect, or if $B(T)=\text{False}$ (indicating the word is not in the dictionary), the system performs a Dictionary Lookup to confirm or correct the word.
Grammar Rules
The system incorporates a set of Zarma grammar rules, a task made challenging by the lack of a standard writing format. This phase scrutinizes elements like consonant rules and vowel lengths, essential for keeping the text accurate to Zarma’s grammatical essence despite its absence of a writing standard.
Correction Algorithm
When an error is detected, either because the Bloom Filter returned false or the grammatical check identified an issue, the Correction Algorithm is activated. This component leverages the Levenshtein distance $L$ to identify the smallest number of edits—insertions, deletions, or substitutions—required to rectify the erroneous word. For illustration, given a corrupted sentence "A sindq biri," the Levenshtein algorithm’s operation can be represented as:
Corrupted Sentence: "A sindq $\begin{array}[]{c}\text{{Corrupted Sentence}: "A sindq biri"}\\ \downarrow\\ \text{{LevSuggest}}("sindq")\\ \downarrow\\ \{\text{"sind"},\text{"sinda"}\}\end{array}$ (3)
Here, "sindq" undergoes comparison with the lexicon, and "sind" and "sinda" are suggested based on their proximity in terms of minimal edit distance. In this context, $L(a,b)$ calculates the distance between the incorrect string $a$ and a potential correct string $b$, considering their lengths $i$ and $j$, with $1_{(a_{i}\neq b_{j})}$ serving as an indicator function to highlight discrepancies. This mechanism ensures the algorithm’s efficiency in providing appropriate corrections, thus enhancing the GEC tool’s overall accuracy for Zarma text.
4 Data Curation Insights
In this section, we detail our approach to gathering and preparing data for training our GEC models for Zarma. We constructed two distinct datasets—a synthetic dataset and a human-annotated “gold” dataset—which we then combined to create a comprehensive training set. The final dataset comprises over 250,000 examples, split into 80% for training, 10% for validation, and 10% for testing.
4.0.1 Synthetic Data
To address the scarcity of annotated data in Zarma, we generated a synthetic dataset by applying a custom corruption script to the Feriji Dataset (Keita et al., 2024). This script introduces realistic typographical and grammatical errors into Zarma sentences, mimicking plausible human mistakes. The corruption process is guided by a defined set of operations to ensure the errors are representative of actual language use.
Specifically, we applied a noise function N to each original sentence SZ (composed of a subject S, verb V, and complement C) to produce a corrupted sentence CSZ:
$CSZ(S,V,C)=N\circ SZ(S,V,C)$ (4)
The noise function N consists of four operations: deletions $\delta$, insertions $\mu$, substitutions $\sigma$, and transpositions $\tau$. These operations are designed to introduce errors without inadvertently creating another grammatically valid Zarma sentence.
For example, consider the original sentence SZ = A go koy fuo” (He is going home”). Applying a substitution operation $\delta$ that changes ’go’ to ’ga’ yields the corrupted sentence CSZ = ’A ga koy fuo’. While ’ga’ is a valid Zarma word, in this context, it alters the meaning to ’He will go home,’ introducing a grammatical error.
$\textbf{SZ}="A\,go\,koy\,fuo"$ (5)
$\textbf{N}(\textbf{SZ})=\textbf{SZ}\xrightarrow{\sigma("go"\to"ga")}"A\,ga\,koy\,fuo"$
$\textbf{CSZ}="A\,ga\,koy\,fuo"$
[…]
Furthermore, to enrich the dataset with diverse linguistic nuances, the script generates four corrupted variants for each correct sentence in the dataset. A snapshot of the synthetic data is provided in Table 1.
Correct Sentence Corrupted Sentences
Sintina gaa Irikoy na beena da ganda taka. Sintina gaa Irikog na beena da ganda taka.
Sintina gaa Irikoy na been da ganda taka.
Sintina aga Irikoy na beena da ganda taka.
Sintina gaa Irikoy na beena ea ganda taka.
Table 1: Snapshot of The Synthetic Data
4.1 Human-Annotated Data
The second dataset, referred to as the ‘gold data,’ was curated by human annotators proficient in Zarma. They introduced grammatical and logical errors (LG)—more details about this errors type can be found in Section B.2—into Zarma sentences and provided justifications for each alteration. This dataset exposes the models to a broader array of error types—particularly LG errors—and their corresponding corrections, enhancing the models’ capacity to generalize and accurately correct unseen texts in Zarma, especially in zero-shot or few-shot learning scenarios. An example from the gold data is shown in Table 2
Incorrect Sentence Correction Error Explanation
Souba, Ay koy Niamey Souba, Ay ga koy Niamey "Souba" means tomorrow, and therefore the tense must be in the future using the future tense marker "ga" after the subject "Ay."
Table 2: Snapshot of the Gold Data
5 Experiment
We conducted a series of experiments to evaluate the effectiveness of different GEC methods for Zarma. We selected three models for training based on their demonstrated proficiency in multilingual tasks and aptitude for few-shot learning: Gemma (Team et al., 2024), MT5-small, and M2M100. The training was conducted on Google Cloud, using a Deep Learning Virtual Machine equipped with NVIDIA GPUs T4 x2. Due to computational resource constraints, we applied QLoRA quantization (Dettmers et al., 2023) to the Gemma model to reduce memory requirements. For MT5-small and M2M100, we used smaller variants of the models to ensure feasible training times. For the LLM-based methods (Gemma and MT5-small), we structured the combined dataset as described in Section 4, incorporating both incorrect sentences and their corrections along with error explanations. For the MT-based method (M2M100), we adopted a MT task-specific format, using aligned pairs of incorrect and correct Zarma sentences without explicit error explanations. The detailed training settings for each model are presented in Table 7.
6 Results & Comparative Analysis
To assess the effectiveness of each GEC method for the types of errors addressed in this paper (as detailed in Section B), we conducted a comprehensive evaluation consisting of two parts: an automatic evaluation (AE) and a manual evaluation (ME) by native Zarma speakers. We included the ME because AE tools are primarily developed and tested on high-resource languages. Therefore, their scores may not reliably reflect performance in low-resource settings like Zarma. The ME allows us to obtain assurances from language speakers regarding the actual effectiveness of the methods.
6.1 Automatic Evaluation Results
In the AE, we employed two sets of metrics to evaluate our techniques. The first set consisted of common machine learning metrics, including the GLEU score (Wang et al., 2018) and the M2 score (Dahlmeier and Ng, 2012). However, these metrics are not applicable to the rule-based method due to its deterministic nature and the way these metrics are designed, which favor probabilistic models that produce varying outputs. The second set focused on spelling correction metrics—measuring the detection rate and suggestion accuracy. The results are presented in Table 3.
Methods AE Evaluation
Common ML Metrics Spelling Correction
GLEU Detection Correction
Rule-based - - 100% 96.27%
Gemma 2b 0.4572 0.528 76.19% 43.28%
MT5-small 0.6203 0.731 90.62% 57.15%
M2M100 0.7896 0.914 95.82% 78.90%
Table 3: AE metrics scores
6.1.1 Common Machine Learning Metrics
The GLEU and M2 scores provide an overall assessment of the models’ performance in correcting grammatical errors by comparing the corrected sentences to reference sentences. The higher scores achieved by the M2M100 model (GLEU: 0.7896, M2: 0.914) indicate its higher performance among the tested models. The LLM-based models, Gemma 2b and MT5-small, obtained moderate scores, reflecting their ability to correct some errors but not as effectively as the M2M100 model.
6.1.2 Spelling Correction Metrics
For the spelling correction evaluation, we used a set of 3,539 sentences from the Feriji Dataset, introducing typographical errors using our corruption script (as described in Section 4). We measured two key metrics:
  • Detection Rate: The percentage of errors that the method successfully identified and attempted to correct.
  • Suggestion Accuracy: The percentage of corrections suggested by the method that were accurate.
As shown in Table 3, the rule-based method achieved a detection rate of 100% and a suggestion accuracy of 96.27%. This impressive performance is primarily due to its extensive dictionary and predefined rules, which allow it to effectively flag and correct errors covered within its scope. The M2M100 model also performed well, with a detection rate of 95.82% and a suggestion accuracy of 78.90%. This suggests that the MT-based approach is capable of effectively identifying and correcting spelling errors, even in a low-resource language setting. The LLM-based models, Gemma 2b and MT5-small, showed lower detection rates and suggestion accuracies. Gemma 2b detected 76.19% of errors with a suggestion accuracy of 43.28%, while MT5-small detected 90.62% of errors with a suggestion accuracy of 57.15%. These results indicate that while LLMs have some capacity to correct spelling errors, they are less effective than the MT-based approach and the rule-based method in this specific task.
6.2 Manual Evaluation Results
Given the limitations of AE metrics in low-resource settings, we conducted the ME involving five native Zarma speakers. The evaluators assessed how well each method corrected LG errors, using a scale from 1 to 5, where 1 indicates a poor correction and 5 indicates an excellent correction. The evaluation was performed on a set of 300 sentences from the Feriji Dataset, which were manually corrupted by our annotators to include LG errors. Detailed information about the evaluation protocol and the recruitment of evaluators can be found in Section C. The results are presented in Table 4.
Methods ME avg (1-5)
Logical Errors Correction Sentence Improvement
Rule-based 0.4 0
Gemma 2b 1 0
MT5-small 1.7 1
M2M100 3 2.5
Table 4: Manual Evaluation
The rule-based method struggled with LG errors, achieving low scores (0.4 for logical error correction and 0.0 for sentence improvement). This outcome is expected, as the rule-based approach relies on predefined rules and lacks the ability to understand the context or semantics of sentences. The M2M100 model outperformed the other methods, achieving higher scores in both logical error correction (3.0) and sentence improvement (2.5). This demonstrates that the MT-based approach is more effective at handling LG errors and improving the overall quality of the sentences. The LLM-based models achieved moderate scores, with MT5-small performing better than Gemma 2b. MT5-small received scores of 1.7 for logical error correction and 1.0 for sentence improvement, suggesting some capability in handling LG errors, but not as effectively as the M2M100 model. We observed that the models faced varying levels of difficulty with different types of LG errors. For instance, they were generally better at correcting verb tense errors than subject-verb agreement errors.
6.3 Zero-Shot scores
To further assess the generalization capabilities of our models, we conducted a zero-shot evaluation using a separate dataset containing sentences with words and grammatical structures not present in the training data. This evaluation is crucial for understanding how well the models can handle new or unseen data, which is a common scenario in low-resource language settings. The results of the zero-shot evaluation are presented in Table 5.
Methods AE Evaluation Manual Evaluation avg (1-5)
Common ML Metrics Spelling Correction Logical Errors Correction Sentence Improvement
GLEU Detection Correction
Rule-based - - 94.17% 81.04% 0 0
Gemma 2b 0.3926 0.481 53.93% 26.52% 0.6 0
MT5-small 0.5018 0.657 73.64% 58.31% 0.8 1
M2M100 0.7683 0.904 92.27% 78.34% 2.7 1.3
Table 5: Performance Evaluation on Zero-Shot Dataset
As expected, the rule-based method showed a decrease in performance compared to the controlled evaluation. Its detection rate dropped to 94.17% and suggestion accuracy to 81.04%. This decline is due to the method’s reliance on a predefined dictionary and rules, which may not cover new words or grammatical constructions present in the zero-shot dataset. Additionally, the ME scored it with 0 points for both logical errors correction and sentence improvement.
The M2M100 model again demonstrated superior performance among the tested models in the zero-shot setting. It achieved a GLEU score of 0.7683 and an M² score of 0.904, indicating strong overall correction capability. Its detection rate was 92.27% with a suggestion accuracy of 78.34%, showing that it can effectively identify and correct errors in unseen data. Furthermore, in ME, it received 2.7 for logical errors correction and 1.3 for sentence improvement.
The LLM-based models, Gemma 2b and MT5-small, showed reduced performance in the zero-shot evaluation. Gemma 2b’s detection rate was 53.93% with a suggestion accuracy of 26.52%, while MT5-small achieved a detection rate of 73.64% and a suggestion accuracy of 58.31%. In terms of ME, Gemma 2b scored 0.6 for logical errors correction and 0 for sentence improvement, while MT5-small achieved 0.8 and 1, respectively.
Our results strongly suggest that MT models trained on extensive multilingual datasets hold significant potential for improving GEC in low-resource languages. The ME scores further validate that the M2M100 model excels not only in automated metrics but also in human judgments, particularly in complex correction tasks. To validate the reproducibility and robustness of our methods, we conducted additional experiments with Bambara—a language from a different linguistic family. The results of this experiment are detailed in Section A.
7 Strengths and Limitations of Each Method
7.1 Rule-Based Method
Rule-based approaches demonstrated high effectiveness in addressing predictable and previously encountered error patterns. In controlled tests, these methods achieved perfect detection and high suggestion accuracy. Their strength lies in the reliance on a comprehensive set of predefined rules and a detailed lexicon. However, this reliance also presents significant limitations—inflexibility and lack of adaptability. Rule-based methods struggle to handle new or unexpected errors, especially in dynamic language use where novel words or grammatical constructions are common. This limitation is particularly pronounced in zero-shot scenarios, where the system encounters words or structures not included in its rules or dictionary. Moreover, the dependency on extensive and carefully curated linguistic resources restricts the scalability of rule-based methods, especially for low-resource languages like Zarma, where such resources are often limited or incomplete (Scannell, 2007).
7.2 LLMs Method
The LLMs in our experiments—Gemma 2b and MT5-small—demonstrated moderate performance in both controlled and zero-shot scenarios. A key strength of LLMs is their capacity to understand context and generate corrections based on broader textual cues, rather than relying only on direct matches to known errors. However, their performance is significantly dependent on the diversity and quality of the training data. A major limitation is that most pre-existing LLMs are primarily trained on data from high-resource languages—mainly Western languages. Consequently, their applicability to African languages like Zarma is hindered by a lack of representative training examples (Bender et al., 2021). This results in lower suggestion accuracy and difficulties in effectively handling the linguistic complexities of these languages. Furthermore, training and fine-tuning LLMs require substantial computational resources, creating a significant barrier in resource-constrained environments.
7.3 MT Method
The MT approach—using the M2M100 model—demonstrated exceptional performance, particularly in zero-shot scenarios, surpassing both rule-based methods and LLMs. The strength of this approach lies in the ability of MT models to leverage multilingual translation mechanisms, effectively adapting to the nuances of diverse languages through exposure to vast and varied training datasets. This characteristic makes MT models particularly suitable for GEC in low-resource languages, as they can infer corrections from patterns learned across multiple languages. This aligns with research highlighting the effectiveness of MT models for cross-lingual transfer learning in various NLP tasks (Conneau et al., 2018). However, a significant challenge in using MT models for GEC in low-resource languages is the frequent scarcity of quality parallel corpora for training. Without sufficient data, the models may generate less accurate or contextually inappropriate corrections (Tiedemann, 2020). Additionally, despite their strengths, MT models require fine-tuning and continuous updating to maintain their accuracy and relevance, as language use evolves.
8 Potential Applications
Our team visited Niamey to present our work to the local Zarma community and gather their feedback. The discussions provided valuable insights into potential applications for our GEC tool and broader language models, highlighting their potential impact on the community.
Content Creation and Educational Resources
One significant comment we received was the potential use of our model to translate coding content and general educational materials for enthusiasts and students. There is a growing interest in technology and programming within the community, but a considerable language barrier exists due to the lack of educational resources in Zarma. By translating and correcting coding tutorials, textbooks, and other educational materials into Zarma, our model can help overcome this challenge. Additionally, the GEC tool can be used to improve general educational content in Zarma, including textbooks, instructional materials, and online courses across various subjects.
Cultural Preservation and Documentation
Feedback from the community highlighted the importance of maintaining accurate written records of folklore, oral histories, and traditional knowledge. The GEC tool can support these efforts by providing a reliable means for transcribing and publishing grammatically accurate texts.
9 Conclusion and Future Work
In this study, we have demonstrated the potential of leveraging LLMs and MT models to address the critical need for effective grammatical error correction GEC tools in low-resource languages, with a focus on Zarma. The challenges inherent to developing GEC tools for Zarma—such as data scarcity and lack of standardized orthography—necessitated innovative approaches beyond those used in high-resource language contexts. To overcome these challenges, we employed a novel methodology combining three key components: (1) the development of a custom corruption script to generate synthetic training data, thereby mitigating the limited availability of annotated Zarma text; (2) the creation of a human-annotated "Gold Data" set, incorporating expert knowledge of Zarma grammar to provide a valuable benchmark for evaluating model performance on complex errors; and (3) the adaptation and fine-tuning of advanced LLMs and MT models—Gemma 2b and M2M100—for the task of Zarma GEC.
This research establishes a baseline for future work in developing GEC tools for Zarma and other low-resource languages. Future research directions include:
Hybrid Approaches
Exploring hybrid models that combine the precision of rule-based methods with the adaptability of LLMs and the robustness of MT models could yield more effective GEC tools.
Data Augmentation
Further expanding and diversifying the training dataset, possibly through community-driven data collection and annotation efforts, to enhance model performance and generalization capabilities.
Continuous Learning Mechanisms
Incorporating mechanisms for continuous learning and user feedback could enable the models to adapt to evolving language use, idiomatic expressions, and regional dialects.
Resource Optimization
Investigating methods to reduce the computational requirements of advanced models, such as model compression or distillation techniques, to make GEC tools more accessible in resource-constrained environments.
Cross-Lingual Transfer Learning
Applying the insights gained from this study to other low-resource languages, leveraging cross-lingual transfer learning to accelerate the development of GEC tools across diverse linguistic contexts.
Appendix A Further Experiment with other Languages
After obtaining promising results for Zarma using LLM and MT-based approaches, we conducted further experiments to validate the reproducibility of these methods—using the M2M100 and MT5 models. We selected the Bambara language for this experiment because it belongs to a different linguistic family, allowing us to evaluate the performance of the approaches on a language outside the Nilo-Saharan family. We used the Bayelemabaga dataset (Vydrin et al., 2022) for Bambara. The same data preparation process described in the methodology section was followed; however, we excluded any human-annotated data to focus only on spelling correction performance. The results are presented in Table 6.
Methods AE Evaluation
Common ML Metrics Spelling Correction
GLEU Detection Correction
MT5-small 0.5713 0.643 87.45% 52.91%
M2M100 0.7293 0.851 94.64% 68.18%
Table 6: Performance Metrics for Bambara
The Bambara experiment demonstrated that the MT-based approach outperformed the LLMs-based one regarding word-level correction metrics. The MT-based approach achieved a detection rate of 94.64% and a suggestion accuracy of 68.18%. In contrast, the LLMs-based approach detected 87.45% of errors and suggested corrections with 52.91% accuracy. The promising results from the Bambara experiment highlight the potential of both LLMs and MT models to improve GEC for low-resource languages significantly. However, they also emphasize the necessity for continued expanding and diversifying training datasets.
Appendix B Errors Being Addressed
In this section, we explain the types of errors GEC methods address. We categorize the errors into two main types: word-level correction (spellchecking) and context-level correction. The context-level correction is further divided into grammar errors, logical errors, and sentence improvement. Below, we define each error type and provide examples to illustrate them.
B.1 Word-Level Correction (Spellchecking)
Word-Level correction involves identifying and correcting typographical errors in individual words. These errors are usually due to misspellings, incorrect usage of characters, or typographical mistakes.
  • Example:
    • Incorrect: Sintina gaa Irikog na beena da ganda taka.
    • Correct: Sintina gaa Irikoy na beena da ganda taka.
B.2 Context-Level Correction
Context-level correction involves errors that go beyond individual words and affect the overall structure and meaning of the sentence. We categorize these errors into logical errors and sentence improvement.
B.2.1 Logical Errors
Logical errors include incorrect verb conjugations, subject-verb agreement issues, incorrect use of grammatical markers, and logical inconsistencies within the sentence. These errors affect the grammatical correctness and logical coherence of the sentence.
  • Example:
    • Incorrect: Souba, Ay koy Niamey. (The time indicator "Souba" means "tomorrow," but the verb "koy" indicates present tense.)
    • Correct: Souba, Ay ga koy Niamey. (The future tense marker "ga" matches the time indicator "Souba.")
B.2.2 Sentence Improvement
Sentence improvement involves enhancing the quality of the sentence by making it more precise, concise, or stylistically appropriate. This category addresses grammatically correct sentences that can be improved for better readability or style.
  • Example:
    • Original: I girbi honkuna i tun be.
    • Improved: I ga girbi suba.
Appendix C Evaluators Recruitment Process
To ensure a thorough and culturally appropriate manual evaluation of our GEC methods, we recruited five native Zarma speakers to serve as evaluators. The primary criteria for selection was their fluency in Zarma—their ability to read and write proficiently in the language. The evaluators were selected through community outreach in Niamey, Niger, leveraging connections with local educational institutions. All selected evaluators had completed secondary education or higher.
Prior to the evaluation, we conducted a training workshop to familiarize the evaluators with the objectives of the study, the evaluation protocols, and the scoring guidelines. During the evaluation process, each evaluator independently assessed a set of 300 sentences that had been manually corrupted to include logical and grammatical errors. They reviewed the corrections made by each GEC method and assigned scores based on accuracy and fluency, where:
  • 1 indicates a poor correction with significant errors remaining or introduced.
  • 5 indicates an excellent correction with all errors appropriately addressed and the sentence reading naturally.
To ensure the reliability of the evaluations, we measured inter-annotator agreement using Fleiss’ kappa coefficient. The resulting kappa value was 0.74, indicating substantial agreement among the evaluators. This suggests that the scoring guidelines were clear and that the evaluators applied them consistently.
All evaluators provided informed consent to participate in the study and were compensated for their time and contributions.
Appendix D Training Details
Models Parameters Training Details
QLoRA GPU Used Lr Loss
Gemma 2b 2 billion Applied NVIDIA P100 $2\times 10^{-4}$ 1.2613
MT5-small 300 million Not Applied NVIDIA T4x2 $2\times 10^{-5}$ 0.0345
M2M100 418 million Not Applied NVIDIA P100 $2\times 10^{-5}$ 0.0214
Table 7: Training Settings for the models
Figure 2: Images of the different GEC tool interfaces. The rule-based on the left and the other approaches on the right

### v2

@@
 Abstract
-Grammatical error correction (GEC) is important for improving written materials for low-resource languages like Zarma—spoken by over 5 million people in West Africa.
-Yet it remains a challenging problem.
-This study compares rule-based methods, machine translation (MT) models, and large language models (LLMs) for GEC in Zarma.
-We evaluate each approach’s effectiveness on our manually-built dataset of over 250,000 examples using synthetic and human-annotated data.
-Our experiments show that the MT-based approach using the M2M100 model outperforms others, achieving a detection rate of 95.82% and a suggestion accuracy of 78.90% in automatic evaluations, and scoring 3.0 out of 5.0 in logical/grammar error correction during MEs by native speakers.
-The rule-based method achieved perfect detection (100%) and high suggestion accuracy (96.27%) for spelling corrections but struggled with context-level errors.
-LLMs like MT5-small showed moderate performance with a detection rate of 90.62% and a suggestion accuracy of 57.15%.
-Our work highlights the potential of MT models to enhance GEC in low-resource languages, paving the way for more inclusive NLP tools.
+Grammatical error correction (GEC) aims to improve quality and readability of texts through accurate correction of linguistic mistakes.
+Previous work has focused on high-resource languages, while low-resource languages lack robust tools.
+However, low-resource languages often face problems such as: non-standard orthography, limited annotated corpora, and diverse dialects, which slows down the development of GEC tools.
+We present a study on GEC for Zarma, spoken by over five million in West Africa.
+We compare three approaches: rule-based methods, machine translation (MT) models, and large language models (LLMs).
+We evaluated them using a dataset of more than 250,000 examples, including synthetic and human-annotated data.
+Our results showed that the MT-based approach using M2M100 outperforms others, with a detection rate of 95.
+82% and a suggestion accuracy of 78.
+90% in automatic evaluations (AE) and an average score of 3.0 out of 5.0 in manual evaluation (ME) from native speakers for grammar and logical corrections.
+The rule-based method was effective for spelling errors but failed on complex context-level errors.
+LLMs—MT5-small—showed moderate performance.
+Our work supports use of MT models to enhance GEC in low-resource settings, and we validated these results with Bambara, another West African language.
 Grammatical Error Correction for Low-Resource Languages: The Case of Zarma
 Mamadou K.
-KEITA1, Christopher Homan3, Sofianou Abdoulaye Hamani1, Adwoa Bremang2, Marcos Zampieri4, Habibatou Abdoulaye Alfari2, Elysabhete Amadou Ibrahim2, Dennis Owusu2 1NILab, 2Ashesi University, 3Rochester Institute of Technology, 4George Mason University
+KEITA1, Christopher Homan3, Marcos Zampieri4, Adwoa Bremang2, Habibatou Abdoulaye Alfari2, Elysabhete Amadou Ibrahim2, Dennis Owusu2 1PAAIR Institute, 2Ashesi University, 3Rochester Institute of Technology, 4George Mason University
 1 Introduction
-Grammatical error correction (GEC) is an essential task in natural language processing (NLP) that aims to improve the quality and readability of texts by correcting grammatical errors.
-GEC are important for enhancing written materials, significantly impacting educational outcomes, professional opportunities, and access to information.
-This is particularly relevant in under-resourced settings where limited access to academic resources and formal training can amplify language proficiency disparities.
-Furthermore, communities in these settings often rely on local languages for communication and knowledge transmission.
-While GEC tools are well-developed for high-resource languages—especially English—where extensive annotated datasets and standardized writing systems are available (Napoles et al., 2019), GEC presents significant challenges for low-resource languages.
-Zarma, a Nilo-Saharan language spoken by over 5 million people across Niger and neighboring countries (Lewis et al., 2016), exemplifies the difficulties faced by many low-resource languages.
-The lack of standardized orthography and the limited availability of annotated data make it challenging to develop practical GEC tools.
-These challenges are not unique to Zarma but are shared by a broad class of low-resource languages, including many African and indigenous languages worldwide.
-Addressing these challenges requires innovative approaches that can work with minimal data and accommodate non-standardized texts.
-The emergence of large language models (LLMs) gave notice to a new era in NLP, empowering machine learning models to learn and generate human-like text across many languages (Devlin et al., 2018; Brown et al., 2020).
-LLMs exhibit remarkable zero-shot and few-shot learning capabilities (Wan et al., 2023), which are potentially beneficial for low-resource languages.
-These capabilities allow LLMs to perform effectively with minimal data, making them valuable for languages with limited annotated datasets.
-However, the use of LLMs is also challenging in low-resource settings.
-They are primarily trained on data from high-resource languages, which may limit their performance when applied to low-resource languages.
-The lack of representative training data can lead to errors and biases, reducing their effectiveness.
-Additionally, the significant computational resources required to fine-tune and deploy LLMs can be a barrier in resource-constrained environments.
-This study investigates the potential of LLMs and traditional models to improve GEC for Zarma by comparing conventional rule-based methods, machine translation (MT)-based models, and the novel application of LLMs for Zarma GEC.
-To explore the performance of these models on other low-resource languages, we replicate our Zarma GEC experiments with Bambara, a West African language spoken in Mali.
-Through a case study, we highlight the strengths and weaknesses of each method, aiming to bridge the linguistic gap in NLP for low-resource languages like Zarma and Bambara.
-Our research addresses the following questions:
+Grammatical error correction (GEC) is important to improving text quality in different scenarios.
+High-resource languages have well-developed GEC tools (Napoles et al., 2019), but low-resource languages have few resources.
+Zarma, a Nilo-Saharan language spoken widely in West Africa (Lewis et al., 2016), faces these challenges.
+Zarma texts lack standardized orthography and large annotated datasets.
+Developers need approaches that handle minimal resources and non-standard writing.
+Large language models (LLMs) hold promise—as they can handle various languages and tasks (Brown et al., 2020)—but performance on low-resource languages can be limited.
+They rely heavily on data, which is scarce for Zarma.
+We investigate three GEC approaches for Zarma: a rule-based method, a machine translation (MT)-based model, and LLMs.
+We also test with Bambara, another West African language, to confirm generalization.
+This paper addresses two key questions:
 RQ1:
-Do state-of-the-art LLM models outperform conventional rule-based and MT-based models on GEC for Zarma texts?
+Do LLMs outperform traditional rule-based and MT models in Zarma GEC?
 RQ2:
-What are the specific strengths and limitations of each GEC approach in low-resource settings, considering the lack of standardized orthography and limited annotated data?
+What are the strengths and limitations of each approach, given the non-standard orthography and limited data?
+RQ3:
+Does the best-performing approach replicate for another low-resource language—Bambara?
 The main contributions of this paper are as follows:
-  • A comprehensive evaluation of three distinct GEC approaches for Zarma (rule-based, MT-based, and LLM-based).
-Our findings show that the MT-based approach delivered the highest accuracy with a detection rate of 95.82%, a suggestion accuracy of 78.90%, and acceptable performance in zero-shot scenarios.
-  • Reproduction of the experiments with Bambara—a manding family language—to confirm replicability and broaden the study’s scope beyond Zarma.
-  • Development and planned public release (upon acceptance) of models fine-tuned for GEC of the tested languages.
+  • A comparison of three GEC methods for Zarma.
+The MT-based approach—M2M100—provided the best results, with a detection rate of 95.
+82% and a suggestion accuracy of 78.
+90% in automatic evaluations (AE) and an average score of 3.0 out of 5.0 in manual evaluation (ME).
+  • A dataset of more than 250,000 synthetic and human-annotated Zarma examples.
+  • Experiments with Bambara to confirm replicability.
+  • Planned release of trained models to promote future research.
 2 Related Work
@@
 The fine-tuning process involves adapting the pre-trained MT model to treat error correction as a "translation" task, where the source language is the incorrect sentence and the target language is the corrected sentence.
-However, the research also identified a fundamental challenge: the trade-off between precision and recall when training with synthetic data.
-This finding shows the need for further research to optimize these models for low-resource GEC tasks, potentially through improved data augmentation techniques or developing specialized architectures for error correction.
+The complexity of balancing precision and recall on synthetic data remains an open problem.
 3 Proposed Methods
 3.1 GEC with LLMs
-LLMs have significantly advanced NLP, exhibiting capabilities in multitasking, few-shot and multilingual learning.
+LLMs have significantly advanced NLP, showing capabilities in multitasking, single-shot, and multilingual learning.
 These models, extensively pre-trained on diverse datasets, demonstrate a remarkable ability to grasp nuanced aspects of language, reasoning, and context (Brown et al., 2020; Raffel et al., 2020).
-This section outlines our methodology for leveraging LLMs to develop a GEC tool for Zarma.
+This section outlines our methodology for using LLMs to develop a GEC tool for Zarma.
 The proposed GEC tool is designed to function independently of predefined grammar rules or lexicons, using the models’ few-shot learning capabilities for enhanced efficiency in low-resource scenarios.
 3.1.1 Implementation
-We employed two distinct approaches for LLM-based GEC:
+We used two distinct approaches for LLM-based GEC:
 Instruction and Error Explanation Fine-tuning
-This method involves embedding training data within a contextual sentence structure, enhancing the model’s reasoning abilities, and facilitating effective learning from the examples.
-We use the following instruction prompt:
+This method involves embedding training data within a contextual sentence structure. We embed examples in a prompt format:
 Prompt: "Zarma sentence: [Incorrect Sentence],Correct the zarma sentence: [Correct Sentence]Error Causes: : [Error $\begin{array}[]{l}\text{{Prompt}: "Zarma sentence: [{Incorrect Sentence}],}\\ \text{Correct the zarma sentence: [{Correct Sentence}]}\\ \text{{Error Causes: }: [{Error Cause}]."}\end{array}$
-This format, particularly the error explanation component, is important for aiding the model’s comprehension of Zarma’s grammatical contexts and patterns, as demonstrated in previous research (Schick and Schütze, 2021; Wei et al., 2021).
+This format helps the model reason about corrections and reflect on why errors occur, as demonstrated in previous research (Schick and Schütze, 2021; Wei et al., 2021).
 Non-Prompt Fine-tuning Using Aligned Sentences
-This approach involves fine-tuning the model directly on parallel data of incorrect and correct Zarma sentences without explicit prompts.
-This leverages the model’s ability to learn the implicit mapping between incorrect and correct forms.
+We also fine-tuned LLMs on parallel incorrect and correct sentences. This leverages the model’s ability to align corrupted text with corrected forms.
 3.2 GEC with MT Models
-Exploring MT models for GEC represents a promising avenue for addressing GEC challenges in low-resource languages like Zarma.
-MT models, particularly those pre-trained on multilingual datasets, offer strong capabilities for processing text across diverse linguistic frameworks.
+Using MT models for GEC represents a promising avenue for addressing GEC challenges in low-resource languages such as Zarma.
+MT models—particularly those pre-trained on multilingual datasets—offer strong capabilities for processing text across diverse linguistic frameworks.
 3.2.1 Implementation
 We chose the M2M100 model (Fan et al., 2020) for its demonstrated ability to translate between many languages, indicating its potential to capture cross-linguistic patterns relevant to GEC.
-To adapt M2M100 for Zarma GEC, we fine-tuned it on a corrupted corpus designed to reflect common errors in Zarma text.
-This corpus was generated by applying a custom noise script, described in Section 4.
-The script introduces various errors, ensuring the training data effectively represents realistic challenges in real-world Zarma text.
-This corrupted corpus, paired with the original correct sentences, serves as the training data for M2M100, enabling the model to learn the mapping from incorrect to correct Zarma.
+We adapt M2M100 for GEC by training it to translate corrupted Zarma text into corrected versions.
+This approach uses a synthetic corpus generated with a noise script (Section 4).
+The script introduces various errors, ensuring that the training data effectively represent realistic challenges in real-world Zarma text.
+This corrupted corpus, paired with the original correct sentences, serves as the training data for M2M100, allowing the model to learn the mapping from incorrect to correct Zarma.
 Detailed training settings and evaluation metrics for this MT-based GEC approach can be found in Tables 7, 4, and 5.
 3.3 Rule-based GEC for Zarma
-As a baseline, we designed a rule-based GEC process, based on Levenshtein distance and the Bloom filter (Bloom, 1970), for Zarma (Figure 1).
+As a baseline, we built a rule-based system using Levenshtein distance and a Bloom filter (Bloom, 1970), for Zarma (Figure 1).
 Additionally, we implemented a tool and API in Python.
-The tool—the first of its kind for Zarma, to our knowledge—is designed to cater to a wide range of users.
+The tool—the first of its kind for Zarma, to our knowledge—is designed for a wide range of users.
 It offers command-line and graphical user interfaces (GUI) and is much less computationally intensive than the LLM-based approach.
-Moreover, our results show that it provides a competitive spell correction performance compared to the LLMs- and MT-based approaches.
-To ensure the tool’s accessibility, we plan a public release upon acceptance.
+Moreover, our results show that it provides a competitive spell correction performance compared to the LLM- and MT-based approaches.
+To ensure the accessibility of the tool, we plan a public release upon acceptance.
 Figure 1: Rule-Based GEC tool Workflow
@@
 At the core of the system is the trie-based dictionary lookup.
+We built this dictionary using the Feriji dataset dictionary (Keita et al., 2024).
 If the Bloom Filter decision is positive ($B(T)=\text{True}$), the system proceeds to a Grammatical Check to verify if the word is grammatically correct.
@@
 This mechanism ensures the algorithm’s efficiency in providing appropriate corrections, thus enhancing the GEC tool’s overall accuracy for Zarma text.
-4 Data Curation Insights
-In this section, we detail our approach to gathering and preparing data for training our GEC models for Zarma.
-We constructed two distinct datasets—a synthetic dataset and a human-annotated “gold” dataset—which we then combined to create a comprehensive training set.
-The final dataset comprises over 250,000 examples, split into 80% for training, 10% for validation, and 10% for testing.
+4 Data
+We created a combined dataset of over 250,000 examples by merging two distinct datasets—a synthetic dataset and a human-annotated ”gold” dataset.
+We split it into 80% training, 10% validation, and 10% testing.
 4.0.1 Synthetic Data
-To address the scarcity of annotated data in Zarma, we generated a synthetic dataset by applying a custom corruption script to the Feriji Dataset (Keita et al., 2024).
-This script introduces realistic typographical and grammatical errors into Zarma sentences, mimicking plausible human mistakes.
+To address the scarcity of annotated data in Zarma, We applied a noise script to the Feriji dataset (Keita et al., 2024).
+This script introduces realistic typographical and grammatical errors into Zarma sentences, to mimick plausible human mistakes.
 The corruption process is guided by a defined set of operations to ensure the errors are representative of actual language use.
-Specifically, we applied a noise function N to each original sentence SZ (composed of a subject S, verb V, and complement C) to produce a corrupted sentence CSZ:
+We applied a noise function N to each original sentence SZ (composed of a subject S, verb V, and complement C) to produce a corrupted sentence CSZ:
 $CSZ(S,V,C)=N\circ SZ(S,V,C)$ (4)
 The noise function N consists of four operations: deletions $\delta$, insertions $\mu$, substitutions $\sigma$, and transpositions $\tau$.
-These operations are designed to introduce errors without inadvertently creating another grammatically valid Zarma sentence.
 For example, consider the original sentence SZ = A go koy fuo” (He is going home”).
@@
 While ’ga’ is a valid Zarma word, in this context, it alters the meaning to ’He will go home,’ introducing a grammatical error.
-$\textbf{SZ}="A\,go\,koy\,fuo"$ (5)
+$\textbf{SZ}="A\,go\,koy\,fuo"(I\ am\ going\ home)$ (5)
 $\textbf{N}(\textbf{SZ})=\textbf{SZ}\xrightarrow{\sigma("go"\to"ga")}"A\,ga\,koy\,fuo"$
-$\textbf{CSZ}="A\,ga\,koy\,fuo"$
+$\textbf{CSZ}="A\,ga\,koy\,fuo"(I\ will\ go\ home)$
 […]
@@
 4.1 Human-Annotated Data
-The second dataset, referred to as the ‘gold data,’ was curated by human annotators proficient in Zarma.
+The second dataset was curated by human annotators proficient in Zarma.
 They introduced grammatical and logical errors (LG)—more details about this errors type can be found in Section B.2—into Zarma sentences and provided justifications for each alteration.
 This dataset exposes the models to a broader array of error types—particularly LG errors—and their corresponding corrections, enhancing the models’ capacity to generalize and accurately correct unseen texts in Zarma, especially in zero-shot or few-shot learning scenarios.
-An example from the gold data is shown in Table 2
+An example from the human-annotated dataset is shown in Table 2
 Incorrect Sentence Correction Error Explanation
 Souba, Ay koy Niamey Souba, Ay ga koy Niamey "Souba" means tomorrow, and therefore the tense must be in the future using the future tense marker "ga" after the subject "Ay."
-Table 2: Snapshot of the Gold Data
+Table 2: Snapshot of the human-annotated dataset
 5 Experiment
@@
 For MT5-small and M2M100, we used smaller variants of the models to ensure feasible training times.
-For the LLM-based methods (Gemma and MT5-small), we structured the combined dataset as described in Section 4, incorporating both incorrect sentences and their corrections along with error explanations.
+For LLM-based methods (Gemma and MT5-small), we structured the combined dataset as described in Section 4, to incorporate both incorrect sentences and their corrections along with error explanations.
 For the MT-based method (M2M100), we adopted a MT task-specific format, using aligned pairs of incorrect and correct Zarma sentences without explicit error explanations.
@@
 6 Results & Comparative Analysis
-To assess the effectiveness of each GEC method for the types of errors addressed in this paper (as detailed in Section B), we conducted a comprehensive evaluation consisting of two parts: an automatic evaluation (AE) and a manual evaluation (ME) by native Zarma speakers.
+To assess the effectiveness of each GEC method for the types of errors addressed in this paper (as detailed in Section B), we conducted an evaluation consisting of two parts: an automatic evaluation (AE) and a manual evaluation (ME) by native Zarma speakers.
 We included the ME because AE tools are primarily developed and tested on high-resource languages.
-Therefore, their scores may not reliably reflect performance in low-resource settings like Zarma.
+Therefore, their scores may not reliably reflect performance in low-resource settings such as Zarma.
 The ME allows us to obtain assurances from language speakers regarding the actual effectiveness of the methods.
@@
 The higher scores achieved by the M2M100 model (GLEU: 0.7896, M2: 0.914) indicate its higher performance among the tested models.
-The LLM-based models, Gemma 2b and MT5-small, obtained moderate scores, reflecting their ability to correct some errors but not as effectively as the M2M100 model.
+The LLM-based models—Gemma 2b and MT5-small—obtained moderate scores, which shows their ability to correct some errors but not as effectively as the M2M100 model.
 6.1.2 Spelling Correction Metrics
@@
 As shown in Table 3, the rule-based method achieved a detection rate of 100% and a suggestion accuracy of 96.27%.
-This impressive performance is primarily due to its extensive dictionary and predefined rules, which allow it to effectively flag and correct errors covered within its scope.
+This high performance is primarily due to its extensive dictionary and predefined rules, which allow it to effectively flag and correct errors covered within its scope.
 The M2M100 model also performed well, with a detection rate of 95.82% and a suggestion accuracy of 78.90%.
 This suggests that the MT-based approach is capable of effectively identifying and correcting spelling errors, even in a low-resource language setting.
-The LLM-based models, Gemma 2b and MT5-small, showed lower detection rates and suggestion accuracies.
+The LLM-based models—Gemma 2b and MT5-small—showed lower detection rates and suggestion accuracies.
 Gemma 2b detected 76.19% of errors with a suggestion accuracy of 43.28%, while MT5-small detected 90.62% of errors with a suggestion accuracy of 57.15%.
@@
 6.3 Zero-Shot scores
-To further assess the generalization capabilities of our models, we conducted a zero-shot evaluation using a separate dataset containing sentences with words and grammatical structures not present in the training data.
-This evaluation is crucial for understanding how well the models can handle new or unseen data, which is a common scenario in low-resource language settings.
+To further assess the generalization capabilities of our models, we conducted a zero-shot evaluation using a separate dataset.
+For that, We gathered a separate set of Zarma texts from books that had not been used during training.
+We then applied the same corruption script described in Section 4 and asked the annotators to include some LG errors.
+containing sentences with words and grammatical structures not present in the training data.
+This evaluation aimed to understand how well the models can handle new or unseen data, which is a common scenario in low-resource language settings.
 The results of the zero-shot evaluation are presented in Table 5.
@@
 Additionally, the ME scored it with 0 points for both logical errors correction and sentence improvement.
-The M2M100 model again demonstrated superior performance among the tested models in the zero-shot setting.
+The M2M100 model again demonstrated higher performance among the tested models in the zero-shot setting.
 It achieved a GLEU score of 0.7683 and an M² score of 0.904, indicating strong overall correction capability.
@@
 In terms of ME, Gemma 2b scored 0.6 for logical errors correction and 0 for sentence improvement, while MT5-small achieved 0.8 and 1, respectively.
-Our results strongly suggest that MT models trained on extensive multilingual datasets hold significant potential for improving GEC in low-resource languages.
+Our results suggest that MT models trained on multilingual datasets hold significant potential for improving GEC in low-resource languages.
 The ME scores further validate that the M2M100 model excels not only in automated metrics but also in human judgments, particularly in complex correction tasks.
@@
 The results of this experiment are detailed in Section A.
-7 Strengths and Limitations of Each Method
-7.1 Rule-Based Method
-Rule-based approaches demonstrated high effectiveness in addressing predictable and previously encountered error patterns.
-In controlled tests, these methods achieved perfect detection and high suggestion accuracy.
-Their strength lies in the reliance on a comprehensive set of predefined rules and a detailed lexicon.
-However, this reliance also presents significant limitations—inflexibility and lack of adaptability.
-Rule-based methods struggle to handle new or unexpected errors, especially in dynamic language use where novel words or grammatical constructions are common.
-This limitation is particularly pronounced in zero-shot scenarios, where the system encounters words or structures not included in its rules or dictionary.
-Moreover, the dependency on extensive and carefully curated linguistic resources restricts the scalability of rule-based methods, especially for low-resource languages like Zarma, where such resources are often limited or incomplete (Scannell, 2007).
-7.2 LLMs Method
-The LLMs in our experiments—Gemma 2b and MT5-small—demonstrated moderate performance in both controlled and zero-shot scenarios.
-A key strength of LLMs is their capacity to understand context and generate corrections based on broader textual cues, rather than relying only on direct matches to known errors.
-However, their performance is significantly dependent on the diversity and quality of the training data.
-A major limitation is that most pre-existing LLMs are primarily trained on data from high-resource languages—mainly Western languages.
-Consequently, their applicability to African languages like Zarma is hindered by a lack of representative training examples (Bender et al., 2021).
-This results in lower suggestion accuracy and difficulties in effectively handling the linguistic complexities of these languages.
-Furthermore, training and fine-tuning LLMs require substantial computational resources, creating a significant barrier in resource-constrained environments.
-7.3 MT Method
-The MT approach—using the M2M100 model—demonstrated exceptional performance, particularly in zero-shot scenarios, surpassing both rule-based methods and LLMs.
-The strength of this approach lies in the ability of MT models to leverage multilingual translation mechanisms, effectively adapting to the nuances of diverse languages through exposure to vast and varied training datasets.
-This characteristic makes MT models particularly suitable for GEC in low-resource languages, as they can infer corrections from patterns learned across multiple languages.
-This aligns with research highlighting the effectiveness of MT models for cross-lingual transfer learning in various NLP tasks (Conneau et al., 2018).
-However, a significant challenge in using MT models for GEC in low-resource languages is the frequent scarcity of quality parallel corpora for training.
-Without sufficient data, the models may generate less accurate or contextually inappropriate corrections (Tiedemann, 2020).
-Additionally, despite their strengths, MT models require fine-tuning and continuous updating to maintain their accuracy and relevance, as language use evolves.
 7 Potential Applications
 Our team visited Niamey to present our work to the local Zarma community and gather their feedback.
-The discussions provided valuable insights into potential applications for our GEC tool and broader language models, highlighting their potential impact on the community.
+The discussions provided valuable insights into potential applications for our GEC tool and broader language models.
 Content Creation and Educational Resources
@@
 Feedback from the community highlighted the importance of maintaining accurate written records of folklore, oral histories, and traditional knowledge.
-The GEC tool can support these efforts by providing a reliable means for transcribing and publishing grammatically accurate texts.
+The GEC tool can support these efforts by providing a means for transcribing and publishing grammatically accurate texts.
 8 Conclusion and Future Work
-In this study, we have demonstrated the potential of leveraging LLMs and MT models to address the critical need for effective grammatical error correction GEC tools in low-resource languages, with a focus on Zarma.
-The challenges inherent to developing GEC tools for Zarma—such as data scarcity and lack of standardized orthography—necessitated innovative approaches beyond those used in high-resource language contexts.
-To overcome these challenges, we employed a novel methodology combining three key components: (1) the development of a custom corruption script to generate synthetic training data, thereby mitigating the limited availability of annotated Zarma text; (2) the creation of a human-annotated "Gold Data" set, incorporating expert knowledge of Zarma grammar to provide a valuable benchmark for evaluating model performance on complex errors; and (3) the adaptation and fine-tuning of advanced LLMs and MT models—Gemma 2b and M2M100—for the task of Zarma GEC.
-This research establishes a baseline for future work in developing GEC tools for Zarma and other low-resource languages. Future research directions include:
-Hybrid Approaches
-Exploring hybrid models that combine the precision of rule-based methods with the adaptability of LLMs and the robustness of MT models could yield more effective GEC tools.
-Data Augmentation
-Further expanding and diversifying the training dataset, possibly through community-driven data collection and annotation efforts, to enhance model performance and generalization capabilities.
-Continuous Learning Mechanisms
-Incorporating mechanisms for continuous learning and user feedback could enable the models to adapt to evolving language use, idiomatic expressions, and regional dialects.
-Resource Optimization
-Investigating methods to reduce the computational requirements of advanced models, such as model compression or distillation techniques, to make GEC tools more accessible in resource-constrained environments.
-Cross-Lingual Transfer Learning
-Applying the insights gained from this study to other low-resource languages, leveraging cross-lingual transfer learning to accelerate the development of GEC tools across diverse linguistic contexts.
+We presented a comparison of rule-based, LLM-based, and MT-based methods for Zarma GEC.
+The MT-based approach performed best in both controlled and zero-shot scenarios.
+We confirmed these findings by testing with Bambara language.
+This research establishes a foundation for future GEC development in other low-resource languages.
+Future work includes hybrid models that combine rule-based precision with MT and LLM adaptability, data augmentation to improve performance, continuous learning mechanisms, resource optimization, and cross-lingual transfer learning to support a wider range of languages.
+9 Limitations
+9.1 Rule-Based Method
+Rule-based approaches rely on predefined linguistic rules and dictionaries, which limit their adaptability.
+They do not handle new or unseen words well.
+In zero-shot scenarios, where the system encounters words or structures beyond its dictionary, it often fails to correct or detect errors.
+Additionally, the creation and maintenance of linguistic resources is difficult for low-resource languages, which often lack standardized orthography (Scannell, 2007).
+9.2 LLMs Method
+LLMs—Gemma 2b and MT5-small—must rely on diverse training data to capture linguistic nuances.
+Yet, existing LLMs are often trained on data from high-resource languages, which reduces performance in languages like Zarma.
+A lack of training samples in these languages causes the model to produce lower-quality suggestions (Bender et al., 2021).
+In addition, fine-tuning LLMs requires significant computational resources, which can be a challenge in low-resource settings.
+9.3 MT Method
+MT models—M2M100 in our studies—learn from parallel corpora that map incorrect text to correct text.
+In low-resource languages, such parallel data is often lacking or incomplete (Tiedemann, 2020).
+Even though multilingual models can leverage patterns from other languages, the absence of sufficient in-language data can lead to contextually imprecise corrections.
+Moreover, these models require targeted fine-tuning and periodic updates as language use evolves.
 Appendix A Further Experiment with other Languages
 After obtaining promising results for Zarma using LLM and MT-based approaches, we conducted further experiments to validate the reproducibility of these methods—using the M2M100 and MT5 models.
-We selected the Bambara language for this experiment because it belongs to a different linguistic family, allowing us to evaluate the performance of the approaches on a language outside the Nilo-Saharan family.
-We used the Bayelemabaga dataset (Vydrin et al., 2022) for Bambara.
+We selected the Bambara language for this experiment because it belongs to a different linguistic family, which allows us to evaluate the performance of the approaches on a language outside the Nilo-Saharan family.
+We used the Bayelemabaga dataset (Vydrin et al., 2022) for this experiment.
 The same data preparation process described in the methodology section was followed; however, we excluded any human-annotated data to focus only on spelling correction performance.
@@
 In contrast, the LLMs-based approach detected 87.45% of errors and suggested corrections with 52.91% accuracy.
-The promising results from the Bambara experiment highlight the potential of both LLMs and MT models to improve GEC for low-resource languages significantly.
+The results from the Bambara experiment highlight the potential of both LLMs and MT models to improve GEC for low-resource languages significantly.
 However, they also emphasize the necessity for continued expanding and diversifying training datasets.
@@
 B.2.1 Logical Errors
-Logical errors include incorrect verb conjugations, subject-verb agreement issues, incorrect use of grammatical markers, and logical inconsistencies within the sentence.
-These errors affect the grammatical correctness and logical coherence of the sentence.
+Logical errors are inconsistencies in sentence meaning that arise from incorrect verb conjugations, subject-verb agreement, or the misuse of grammatical markers.
+Unlike grammatical errors, logical errors affect the semantic coherence of a sentence.
   • Example:
@@
 B.2.2 Sentence Improvement
-Sentence improvement involves enhancing the quality of the sentence by making it more precise, concise, or stylistically appropriate.
-This category addresses grammatically correct sentences that can be improved for better readability or style.
+Sentence improvement focuses on enhancing readability, precision, and style. While the sentences may already be grammatically correct, improvements ensure better flow and clarity.
   • Example:
@@
 The primary criteria for selection was their fluency in Zarma—their ability to read and write proficiently in the language.
-The evaluators were selected through community outreach in Niamey, Niger, leveraging connections with local educational institutions.
+The evaluators were selected through community outreach in Niamey, Niger.
 All selected evaluators had completed secondary education or higher.
@@
   • 5 indicates an excellent correction with all errors appropriately addressed and the sentence reading naturally.
-To ensure the reliability of the evaluations, we measured inter-annotator agreement using Fleiss’ kappa coefficient.
-The resulting kappa value was 0.74, indicating substantial agreement among the evaluators.
-This suggests that the scoring guidelines were clear and that the evaluators applied them consistently.
 All evaluators provided informed consent to participate in the study and were compensated for their time and contributions.
@@
 Models Parameters Training Details
-QLoRA GPU Used Lr Loss
-Gemma 2b 2 billion Applied NVIDIA P100 $2\times 10^{-4}$ 1.2613
-MT5-small 300 million Not Applied NVIDIA T4x2 $2\times 10^{-5}$ 0.0345
-M2M100 418 million Not Applied NVIDIA P100 $2\times 10^{-5}$ 0.0214
+QLoRA GPU Used Lr
+Gemma 2b 2 billion Applied P100 $2\times 10^{-4}$
+MT5-small 300 million Not Applied T4x2 $2\times 10^{-5}$
+M2M100 418 million Not Applied P100 $2\times 10^{-5}$
 Table 7: Training Settings for the models
+Appendix E Other Figures
 Figure 2: Images of the different GEC tool interfaces. The rule-based on the left and the other approaches on the right

### v3

@@
 Abstract
-Grammatical error correction (GEC) aims to improve quality and readability of texts through accurate correction of linguistic mistakes.
-Previous work has focused on high-resource languages, while low-resource languages lack robust tools.
-However, low-resource languages often face problems such as: non-standard orthography, limited annotated corpora, and diverse dialects, which slows down the development of GEC tools.
-We present a study on GEC for Zarma, spoken by over five million in West Africa.
+Grammatical error correction (GEC) aims to improve text quality and readability.
+Previous work on the task focused primarily on high-resource languages, while low-resource languages lack robust tools.
+To address this shortcoming, we present a study on GEC for Zarma, a language spoken by over five million people in West Africa.
 We compare three approaches: rule-based methods, machine translation (MT) models, and large language models (LLMs).
-We evaluated them using a dataset of more than 250,000 examples, including synthetic and human-annotated data.
-Our results showed that the MT-based approach using M2M100 outperforms others, with a detection rate of 95.
-82% and a suggestion accuracy of 78.
-90% in automatic evaluations (AE) and an average score of 3.0 out of 5.0 in manual evaluation (ME) from native speakers for grammar and logical corrections.
+We evaluated GEC models using a dataset of more than 250,000 examples, including synthetic and human-annotated data.
+Our results showed that the MT-based approach using M2M100 outperforms others, with a detection rate of 95.82% and a suggestion accuracy of 78.90% in automatic evaluations (AE) and an average score of 3.0 out of 5.0 in manual evaluation (ME) from native speakers for grammar and logical corrections.
 The rule-based method was effective for spelling errors but failed on complex context-level errors.
-LLMs—MT5-small—showed moderate performance.
+LLMs—Gemma 2b and MT5-small—showed moderate performance.
 Our work supports use of MT models to enhance GEC in low-resource settings, and we validated these results with Bambara, another West African language.
 Grammatical Error Correction for Low-Resource Languages: The Case of Zarma
-Mamadou K.
-KEITA1, Christopher Homan3, Marcos Zampieri4, Adwoa Bremang2, Habibatou Abdoulaye Alfari2, Elysabhete Amadou Ibrahim2, Dennis Owusu2 1PAAIR Institute, 2Ashesi University, 3Rochester Institute of Technology, 4George Mason University
+Mamadou K. Keita1, Adwoa Bremang2, Huy Le1 Dennis Owusu2, Marcos Zampieri3, Christopher Homan1 1Rochester Institute of Technology, 2Ashesi University, 3George Mason University
 1 Introduction
 Grammatical error correction (GEC) is important to improving text quality in different scenarios.
-High-resource languages have well-developed GEC tools (Napoles et al., 2019), but low-resource languages have few resources.
-Zarma, a Nilo-Saharan language spoken widely in West Africa (Lewis et al., 2016), faces these challenges.
-Zarma texts lack standardized orthography and large annotated datasets.
-Developers need approaches that handle minimal resources and non-standard writing.
-Large language models (LLMs) hold promise—as they can handle various languages and tasks (Brown et al., 2020)—but performance on low-resource languages can be limited.
-They rely heavily on data, which is scarce for Zarma.
-We investigate three GEC approaches for Zarma: a rule-based method, a machine translation (MT)-based model, and LLMs.
+High-resource languages have well-developed GEC tools (Bryant et al., 2023; Ng et al., 2025) while low-resource languages have few resources.
+Zarma—a Nilo-Saharan language spoken by over five million people, primarily in Niger, where it is a major national language, as well as in Nigeria, Mali, and Burkina Faso (Lewis et al., 2016)—faces these challenges.
+Zarma lacks a standardized orthography and large annotated datasets.
+To create language technology for low-resource languages like Zara, developers need approaches that handle minimal resources and non-standard writing Tapo et al.
+(2020, 2025).
+Large language models (LLMs) hold promise as they are capable of handling various languages and tasks (Brown et al., 2020), but performance on very low-resource languages can be limited.
+LLMs rely heavily on data, which is scarce for Zarma.
+To address these important challenges in low-resource NLP, we investigate GEC approaches for Zarma.
+We evaluate three approaches: a rule-based method, a machine translation (MT)-based model, and LLMs.
 We also test with Bambara, another West African language, to confirm generalization.
-This paper addresses two key questions:
+This paper addresses three key research questions (RQs):
 RQ1:
@@
 The main contributions of this paper are as follows:
+  1. 1.
     A comparison of three GEC methods for Zarma.
-The MT-based approach—M2M100—provided the best results, with a detection rate of 95.
-82% and a suggestion accuracy of 78.
-90% in automatic evaluations (AE) and an average score of 3.0 out of 5.0 in manual evaluation (ME).
-  • A dataset of more than 250,000 synthetic and human-annotated Zarma examples.
+The MT-based approach—M2M100—provided the best results, with a detection rate of 95.82% and a suggestion accuracy of 78.90% in automatic evaluations (AE) and an average score of 3.0 out of 5.0 in manual evaluation (ME).
+  2. 2.
+    A dataset of 250,000 synthetic and human-annotated Zarma examples  [^1] The latest version of the dataset can be found here: 27Group/noisy_zarma.
+  3. 3.
     Experiments with Bambara to confirm replicability.
-  • Planned release of trained models to promote future research.
 2 Related Work
@@
 Another promising approach to GEC leverages pre-trained multilingual MT models.
-The study by (Luhtaru et al., 2024) introduced the "No Error Left Behind" approach, which uses models like MT5 and NLLB (Team et al., 2022).
-The fine-tuning process involves adapting the pre-trained MT model to treat error correction as a "translation" task, where the source language is the incorrect sentence and the target language is the corrected sentence.
+The study by (Luhtaru et al., 2024) introduced the No Error Left Behind approach, which uses models like MT5 and NLLB (Team et al., 2022).
+The fine-tuning process involves adapting the pre-trained MT model to treat error correction as a translation task, where the source language is the incorrect sentence and the target language is the corrected sentence.
 The complexity of balancing precision and recall on synthetic data remains an open problem.
-3 Proposed Methods
+3 Methods
 3.1 GEC with LLMs
 LLMs have significantly advanced NLP, showing capabilities in multitasking, single-shot, and multilingual learning.
-These models, extensively pre-trained on diverse datasets, demonstrate a remarkable ability to grasp nuanced aspects of language, reasoning, and context (Brown et al., 2020; Raffel et al., 2020).
+These models, extensively pre-trained on diverse datasets, demonstrate a remarkable ability to grasp nuanced aspects of language, reasoning, and context (Brown et al., 2020; Raffel and others, 2020).
 This section outlines our methodology for using LLMs to develop a GEC tool for Zarma.
-The proposed GEC tool is designed to function independently of predefined grammar rules or lexicons, using the models’ few-shot learning capabilities for enhanced efficiency in low-resource scenarios.
+The proposed GEC tool is designed to function independently of predefined grammar rules or lexicons, using the models' few-shot learning capabilities for enhanced efficiency in low-resource scenarios.
 3.1.1 Implementation
@@
 This method involves embedding training data within a contextual sentence structure. We embed examples in a prompt format:
-Prompt: "Zarma sentence: [Incorrect Sentence],Correct the zarma sentence: [Correct Sentence]Error Causes: : [Error $\begin{array}[]{l}\text{{Prompt}: "Zarma sentence: [{Incorrect Sentence}],}\\ \text{Correct the zarma sentence: [{Correct Sentence}]}\\ \text{{Error Causes: }: [{Error Cause}]."}\end{array}$
-This format helps the model reason about corrections and reflect on why errors occur, as demonstrated in previous research (Schick and Schütze, 2021; Wei et al., 2021).
+[…]
+This format helps the model reason about corrections and reflect on why errors occur, as demonstrated in previous research (Schick and Schütze, 2021; Wei and others, 2021).
 Non-Prompt Fine-tuning Using Aligned Sentences
-We also fine-tuned LLMs on parallel incorrect and correct sentences. This leverages the model’s ability to align corrupted text with corrected forms.
+We also fine-tuned LLMs on parallel incorrect and correct sentences. This leverages the model's ability to align corrupted text with corrected forms.
 3.2 GEC with MT Models
@@
 This corrupted corpus, paired with the original correct sentences, serves as the training data for M2M100, allowing the model to learn the mapping from incorrect to correct Zarma.
-Detailed training settings and evaluation metrics for this MT-based GEC approach can be found in Tables 7, 4, and 5.
+Detailed training settings and evaluation metrics for this MT-based GEC approach can be found in Tables 8, 5, and 6.
 3.3 Rule-based GEC for Zarma
@@
 The system begins by processing the text, where the Zarma content is segmented into words, punctuation, and spaces using regular expressions.
-This initial step ensures that every part of the text is ready for individual examination—a necessity given Zarma’s linguistic intricacies.
+This initial step ensures that every part of the text is ready for individual examination—a necessity given Zarma's linguistic intricacies.
 Bloom Filter
@@
 The system incorporates a set of Zarma grammar rules, a task made challenging by the lack of a standard writing format.
-This phase scrutinizes elements like consonant rules and vowel lengths, essential for keeping the text accurate to Zarma’s grammatical essence despite its absence of a writing standard.
+Zarma is classified as an agglutinative language, meaning it depends heavily on attaching suffixes to root words to indicate grammatical features like plurality and definiteness (e.g., the definite plural suffix -ey).
+This morphology imposes specific constraints on our rule-based approach: the rules must be robust enough to parse and validate these suffix chains rather than treating words as static units.
+This phase scrutinizes elements like consonant rules and vowel lengths, essential for keeping the text accurate to Zarma's grammatical essence despite its absence of a writing standard.
 Correction Algorithm
@@
 This component leverages the Levenshtein distance $L$ to identify the smallest number of edits—insertions, deletions, or substitutions—required to rectify the erroneous word.
-For illustration, given a corrupted sentence "A sindq biri," the Levenshtein algorithm’s operation can be represented as:
-Corrupted Sentence: "A sindq $\begin{array}[]{c}\text{{Corrupted Sentence}: "A sindq biri"}\\ \downarrow\\ \text{{LevSuggest}}("sindq")\\ \downarrow\\ \{\text{"sind"},\text{"sinda"}\}\end{array}$ (3)
-Here, "sindq" undergoes comparison with the lexicon, and "sind" and "sinda" are suggested based on their proximity in terms of minimal edit distance.
+For illustration, given a corrupted sentence ``A sindq biri,'' the Levenshtein algorithm's operation can be represented as:
+$\begin{array}[]{c}\text{{Corrupted Sentence}: ``A sindq biri''}\\ \downarrow\\ \text{{LevSuggest}}(``sindq^{\prime\prime})\\ \downarrow\\ \{\text{``sind''},\text{``sinda''}\}\end{array}$ (3)
+Here, ``sindq'' undergoes comparison with the lexicon, and ``sind'' and ``sinda'' are suggested based on their proximity in terms of minimal edit distance.
 In this context, $L(a,b)$ calculates the distance between the incorrect string $a$ and a potential correct string $b$, considering their lengths $i$ and $j$, with $1_{(a_{i}\neq b_{j})}$ serving as an indicator function to highlight discrepancies.
-This mechanism ensures the algorithm’s efficiency in providing appropriate corrections, thus enhancing the GEC tool’s overall accuracy for Zarma text.
+This mechanism ensures the algorithm's efficiency in providing appropriate corrections, thus enhancing the GEC tool's overall accuracy for Zarma text.
 4 Data
-We created a combined dataset of over 250,000 examples by merging two distinct datasets—a synthetic dataset and a human-annotated ”gold” dataset.
-We split it into 80% training, 10% validation, and 10% testing.
-4.0.1 Synthetic Data
-To address the scarcity of annotated data in Zarma, We applied a noise script to the Feriji dataset (Keita et al., 2024).
-This script introduces realistic typographical and grammatical errors into Zarma sentences, to mimick plausible human mistakes.
-The corruption process is guided by a defined set of operations to ensure the errors are representative of actual language use.
-We applied a noise function N to each original sentence SZ (composed of a subject S, verb V, and complement C) to produce a corrupted sentence CSZ:
-$CSZ(S,V,C)=N\circ SZ(S,V,C)$ (4)
-The noise function N consists of four operations: deletions $\delta$, insertions $\mu$, substitutions $\sigma$, and transpositions $\tau$.
-For example, consider the original sentence SZ = A go koy fuo” (He is going home”).
-Applying a substitution operation $\delta$ that changes ’go’ to ’ga’ yields the corrupted sentence CSZ = ’A ga koy fuo’.
-While ’ga’ is a valid Zarma word, in this context, it alters the meaning to ’He will go home,’ introducing a grammatical error.
-$\textbf{SZ}="A\,go\,koy\,fuo"(I\ am\ going\ home)$ (5)
-$\textbf{N}(\textbf{SZ})=\textbf{SZ}\xrightarrow{\sigma("go"\to"ga")}"A\,ga\,koy\,fuo"$
-$\textbf{CSZ}="A\,ga\,koy\,fuo"(I\ will\ go\ home)$
-[…]
-Furthermore, to enrich the dataset with diverse linguistic nuances, the script generates four corrupted variants for each correct sentence in the dataset.
-A snapshot of the synthetic data is provided in Table 1.
+To address the scarcity of annotated data in Zarma, we created a combined dataset of 250,000 examples by merging two distinct datasets—a synthetic dataset and a human-annotated ``gold'' dataset.
+The synthetic dataset contains 248,000 sentence pairs; the human-annotated dataset includes 2,000 sentence pairs.
+We split this combined dataset into 80% training, 10% validation, and 10% testing sets.
+4.1 Synthetic Data
 Correct Sentence Corrupted Sentences
@@
 Table 1: Snapshot of The Synthetic Data
+We applied a noise script to the Feriji dataset (Keita et al., 2024) to generate synthetic errors.
+This script introduces typographical and grammatical errors into Zarma sentences—errors that mimic human mistakes.
+To ensure these errors reflect real-world patterns, we worked with five native Zarma speakers—the same individuals who later conducted the manual evaluation (ME) as described in Section B.
+We provided them with audio recordings of 46 Zarma sentences from everyday conversations, which they transcribed into written Zarma—with instructions to not attempt to correct any mistake.
+All typed in French at an average of 38 words per minute using an AZERTY keyboard.
+We compared their transcriptions with the correct Zarma text and analyzed the frequency of error types.
+The study revealed common errors: insertions (e.g., ``Irikoy'' to ``Irikoyu''), and deletion (e.g., ``alfarey'' to ``alfary''); with frequencies of 23% and 12%, respectively.
+Minor transpositions and double vowels were noticed too—less than 2% and 7% respectively.
+The AZERTY keyboard layout and typing speed influence error types such as insertions due to adjacent key proximity—e.g, `u' near `y' in `Irikoyu').
+Also, the average error per sentence was 1 error for short sentences (1-5 words), 3 errors for long sentences (5+ words).
+The noise function $\mathbf{N}$ consists of four operations: deletions $\mathbf{\delta}$, insertions $\mathbf{\mu}$, substitutions $\mathbf{\sigma}$, and transpositions $\mathbf{\tau}$.
+For example, applying a substitution operation to the original sentence $\mathbf{SZ}=\textit{A go koy fuo}$ (``He is going home'') might yield $\mathbf{CSZ}=\textit{A ga koy fuo}$ (``He will go home''), introducing a grammatical error by changing the tense.
+$\mathbf{SZ}=``A\,go\,koy\,fuo^{\prime\prime}$ (4)
+$\mathbf{N}(\mathbf{SZ})=\mathbf{SZ}\xrightarrow{\sigma(``go^{\prime\prime}\to``ga^{\prime\prime})}``A\,ga\,koy\,fuo^{\prime\prime}$
+$\mathbf{CSZ}=``A\,ga\,koy\,fuo^{\prime\prime}$
+$\mathbf{SZ}\xrightarrow{N}\mathbf{Correct}\,\mathbf{CSZ}$
+Furthermore, to enrich the dataset with diverse linguistic nuances, the script generates four corrupted variants for each correct sentence in the dataset.
+A snapshot of the synthetic data appears in Table 1.
 4.2 Human-Annotated Data
-The second dataset was curated by human annotators proficient in Zarma.
-They introduced grammatical and logical errors (LG)—more details about this errors type can be found in Section B.2—into Zarma sentences and provided justifications for each alteration.
-This dataset exposes the models to a broader array of error types—particularly LG errors—and their corresponding corrections, enhancing the models’ capacity to generalize and accurately correct unseen texts in Zarma, especially in zero-shot or few-shot learning scenarios.
+The human-annotated dataset was curated by native Zarma speakers who introduced logical and grammatical (LG) errors into Zarma sentences and provided justifications for each alteration.
+This dataset exposes the models to a broader array of error types—particularly logical errors—and their corresponding corrections, enhancing the models' capacity to generalize and accurately correct unseen texts in Zarma.
 An example from the human-annotated dataset is shown in Table 2
 Incorrect Sentence Correction Error Explanation
-Souba, Ay koy Niamey Souba, Ay ga koy Niamey "Souba" means tomorrow, and therefore the tense must be in the future using the future tense marker "ga" after the subject "Ay."
+Souba, Ay koy Niamey Souba, Ay ga koy Niamey ``Souba'' means tomorrow, and therefore the tense must be in the future using the future tense marker ``ga'' after the subject ``Ay.''
 Table 2: Snapshot of the human-annotated dataset
@@
 6 Results & Comparative Analysis
-To assess the effectiveness of each GEC method for the types of errors addressed in this paper (as detailed in Section B), we conducted an evaluation consisting of two parts: an automatic evaluation (AE) and a manual evaluation (ME) by native Zarma speakers.
+To assess the effectiveness of each GEC method for the types of errors addressed in this paper (as detailed in Section A), we conducted an evaluation consisting of two parts: an automatic evaluation (AE) and ME by native Zarma speakers.
 We included the ME because AE tools are primarily developed and tested on high-resource languages.
@@
 The first set consisted of common machine learning metrics, including the GLEU score (Wang et al., 2018) and the M2 score (Dahlmeier and Ng, 2012).
-However, these metrics are not applicable to the rule-based method due to its deterministic nature and the way these metrics are designed, which favor probabilistic models that produce varying outputs.
-The second set focused on spelling correction metrics—measuring the detection rate and suggestion accuracy.
+The second set focused on spelling correction metrics—measuring the detection rate, suggestion accuracy, and false positive rate.
 The results are presented in Table 3.
@@
 Common ML Metrics Spelling Correction
-GLEU Detection Correction
-Rule-based - - 100% 96.27%
-Gemma 2b 0.4572 0.528 76.19% 43.28%
-MT5-small 0.6203 0.731 90.62% 57.15%
-M2M100 0.7896 0.914 95.82% 78.90%
+GLEU Detection Correction False Positive
+Rule-based 0.3124 0.401 100% 96.27% 2.5%
+Gemma 2b 0.4572 0.528 76.19% 43.28% 15.3%
+MT5-small 0.6203 0.731 90.62% 57.15% 8.7%
+M2M100 0.7896 0.914 95.82% 78.90% 4.2%
 Table 3: AE metrics scores
 6.1.1 Common Machine Learning Metrics
-The GLEU and M2 scores provide an overall assessment of the models’ performance in correcting grammatical errors by comparing the corrected sentences to reference sentences.
-The higher scores achieved by the M2M100 model (GLEU: 0.7896, M2: 0.914) indicate its higher performance among the tested models.
+The GLEU and M2 scores provide an overall assessment of the models' performance in correcting grammatical errors by comparing the corrected sentences to reference sentences.
+The rule-based method achieved modest scores (GLEU: 0.3124, M2: 0.401), reflecting its effectiveness in spelling correction but limitations in handling complex grammatical errors.
+The higher scores achieved by the M2M100 model (GLEU: 0.7896, M2: 0.914) indicate its superior performance among the tested models.
 The LLM-based models—Gemma 2b and MT5-small—obtained moderate scores, which shows their ability to correct some errors but not as effectively as the M2M100 model.
@@
 For the spelling correction evaluation, we used a set of 3,539 sentences from the Feriji Dataset, introducing typographical errors using our corruption script (as described in Section 4).
-We measured two key metrics:
+We measured three key metrics:
   • Detection Rate: The percentage of errors that the method successfully identified and attempted to correct.
   • Suggestion Accuracy: The percentage of corrections suggested by the method that were accurate.
-As shown in Table 3, the rule-based method achieved a detection rate of 100% and a suggestion accuracy of 96.27%.
-This high performance is primarily due to its extensive dictionary and predefined rules, which allow it to effectively flag and correct errors covered within its scope.
-The M2M100 model also performed well, with a detection rate of 95.82% and a suggestion accuracy of 78.90%.
-This suggests that the MT-based approach is capable of effectively identifying and correcting spelling errors, even in a low-resource language setting.
-The LLM-based models—Gemma 2b and MT5-small—showed lower detection rates and suggestion accuracies.
-Gemma 2b detected 76.19% of errors with a suggestion accuracy of 43.28%, while MT5-small detected 90.62% of errors with a suggestion accuracy of 57.15%.
-These results indicate that while LLMs have some capacity to correct spelling errors, they are less effective than the MT-based approach and the rule-based method in this specific task.
+  • False Positive Rate [^2] While we did not evaluate the models on entirely correct sentences, our False Positive Rate reflects over-corrections within mixed correct-and-incorrect datasets, approximating real-world GEC scenarios where fully pristine text is rare.: The percentage of correct words incorrectly flagged as errors.
+Here, a false positive is counted not only when an entire sentence is flagged, but also when parts of a sentence that are correct are unnecessarily altered.
+In other words, if a sentence contains both correct and incorrect segments, any modification of the correct segments is considered a false positive.
+As shown in Table 3, the rule-based method achieved a detection rate of 100% and a suggestion accuracy of 96.27%, with a low false positive rate of 2.5%.
+This high performance is primarily due to its dictionary—built from the Feriji Dataset—and predefined rules, which allow it to effectively flag and correct errors within its scope while rarely misidentifying correct words.
+The M2M100 model also performed well, with a detection rate of 95.82%, a suggestion accuracy of 78.90%, and a false positive rate of 4.2%, indicating its capability to identify and correct spelling errors in a with minimal over-correction.
+The LLM-based models—Gemma 2b and MT5-small—showed lower detection rates and suggestion accuracies, coupled with higher false positive rates.
+Gemma 2b detected 76.19% of errors with a suggestion accuracy of 43.28% and a false positive rate of 15.3%, while MT5-small detected 90.62% of errors with a suggestion accuracy of 57.15% and a false positive rate of 8.7%.
+These results suggest that while LLMs can correct some spelling errors, they are less effective than the MT-based and rule-based methods in this task and are more prone to incorrectly flagging correct words.
+The reasons of this poor performance of LLMs is due to the fact they were not pre-trained on Zarma, thereby limiting their accuracies in Zarma related tasks.
+6.2 Performance on Synthetic and Human-Annotated Subsets
+To further explore how each method handles different data sources, we tested performance on the synthetic and human-annotated subsets described in Section 4. Table 4 presents the results side by side.
+Methods Synthetic Subset Human-Annotated Subset
+GLEU GLEU
+Rule-based 0.3150 0.405 0.2900 0.375
+Gemma 2b 0.4601 0.531 0.4213 0.492
+MT5-small 0.6257 0.738 0.5812 0.692
+M2M100 0.7923 0.917 0.7618 0.891
+Table 4: Performance metrics on synthetic and human-annotated subsets
+Table 4 reveals a pattern—all models, including the rule-based method, score lower on the human-annotated subset than on the synthetic subset.
+For the rule-based method, GLEU drops from 0.31 to 0.2900 and M² from 0.40 to 0.375, indicating its reduced capability with complex logical errors present in human-annotated data.
+This difference arises because human-annotated data includes complex logical and context-level errors—challenges beyond the simpler typographical errors in the synthetic set.
+Nevertheless, M2M100 maintains its lead across both subsets.
+Its GLEU drops from 0.7923 to 0.7618 and M² from 0.917 to 0.891, yet these scores still outperform Gemma 2b and MT5-small by 5%–10%.
+This consistency highlights M2M100’s strength in tackling diverse error types.
 6.3 Manual Evaluation Results
@@
 The evaluation was performed on a set of 300 sentences from the Feriji Dataset, which were manually corrupted by our annotators to include LG errors.
-Detailed information about the evaluation protocol and the recruitment of evaluators can be found in Section C.
+Detailed information about the evaluation protocol and the recruitment of evaluators can be found in Section B.
 The results are presented in Table 5.
@@
 6.4 Zero-Shot Scores
+Methods AE Evaluation Manual Evaluation avg (1-5)
+Common ML Metrics Spelling Correction Logical Errors Correction Sentence Improvement
+GLEU Detection Correction False Positive
+Rule-based 0.2800 0.365 94.17% 81.04% 3.8% 0 0
+Gemma 2b 0.3926 0.481 53.93% 26.52% 18.2% 0.6 0
+MT5-small 0.5018 0.657 73.64% 58.31% 10.5% 0.8 1
+M2M100 0.7683 0.904 92.27% 78.34% 5.1% 2.7 1.3
+Table 6: Performance Evaluation on Zero-Shot Dataset
+Methods AE Evaluation
+Common ML Metrics Spelling Correction
+GLEU Detection Correction
+MT5-small 0.5713 0.643 87.45% 52.91%
+M2M100 0.7293 0.851 94.64% 68.18%
+Table 7: Performance Metrics for Bambara
 To further assess the generalization capabilities of our models, we conducted a zero-shot evaluation using a separate dataset.
-For that, We gathered a separate set of Zarma texts from books that had not been used during training.
-We then applied the same corruption script described in Section 4 and asked the annotators to include some LG errors.
-containing sentences with words and grammatical structures not present in the training data.
+For that, we gathered a separate set—150 sentences—of Zarma texts from books that had not been used during training.
+We then applied the same corruption script and asked the annotators to include some LG errors, containing sentences with words and grammatical structures not present in the training data.
 This evaluation aimed to understand how well the models can handle new or unseen data, which is a common scenario in low-resource language settings.
 The results of the zero-shot evaluation are presented in Table 6.
-Methods AE Evaluation Manual Evaluation avg (1-5)
-Common ML Metrics Spelling Correction Logical Errors Correction Sentence Improvement
-GLEU Detection Correction
-Rule-based - - 94.17% 81.04% 0 0
-Gemma 2b 0.3926 0.481 53.93% 26.52% 0.6 0
-MT5-small 0.5018 0.657 73.64% 58.31% 0.8 1
-M2M100 0.7683 0.904 92.27% 78.34% 2.7 1.3
-Table 5: Performance Evaluation on Zero-Shot Dataset
 As expected, the rule-based method showed a decrease in performance compared to the controlled evaluation.
-Its detection rate dropped to 94.17% and suggestion accuracy to 81.04%.
-This decline is due to the method’s reliance on a predefined dictionary and rules, which may not cover new words or grammatical constructions present in the zero-shot dataset.
+Its GLEU score dropped to 0.2800 and M² to 0.365, with a detection rate of 94.17%, suggestion accuracy of 81.04%, and a false positive rate of 3.8%.
+This decline is due to the method's reliance on a predefined dictionary and rules, which may not cover new words or grammatical constructions present in the zero-shot dataset, though it maintains a low rate of incorrect flagging.
 Additionally, the ME scored it with 0 points for both logical errors correction and sentence improvement.
@@
 It achieved a GLEU score of 0.7683 and an M² score of 0.904, indicating strong overall correction capability.
-Its detection rate was 92.27% with a suggestion accuracy of 78.34%, showing that it can effectively identify and correct errors in unseen data.
+Its detection rate was 92.27% with a suggestion accuracy of 78.34% and a false positive rate of 5.1%, showing that it can effectively identify and correct errors in unseen data while keeping over-correction low.
 Furthermore, in ME, it received 2.7 for logical errors correction and 1.3 for sentence improvement.
-The LLM-based models, Gemma 2b and MT5-small, showed reduced performance in the zero-shot evaluation.
-Gemma 2b’s detection rate was 53.93% with a suggestion accuracy of 26.52%, while MT5-small achieved a detection rate of 73.64% and a suggestion accuracy of 58.31%.
+The LLM-based models, showed reduced performance in the zero-shot evaluation.
+Gemma 2b's detection rate was 53.93% with a suggestion accuracy of 26.52% and a false positive rate of 18.2%, while MT5-small achieved a detection rate of 73.64%, a suggestion accuracy of 58.31%, and a false positive rate of 10.5%.
 In terms of ME, Gemma 2b scored 0.6 for logical errors correction and 0 for sentence improvement, while MT5-small achieved 0.8 and 1, respectively.
+These higher false positive rates indicate a greater tendency to misidentify correct words as errors in unseen data.
 Our results suggest that MT models trained on multilingual datasets hold significant potential for improving GEC in low-resource languages.
 The ME scores further validate that the M2M100 model excels not only in automated metrics but also in human judgments, particularly in complex correction tasks.
-To validate the reproducibility and robustness of our methods, we conducted additional experiments with Bambara—a language from a different linguistic family.
-The results of this experiment are detailed in Section A.
-7 Potential Applications
+The lower accuracy of LLMs likely arises from their lack of initial pre-training on Zarma and the use of quantization during training, which may impact performance due to limited computational resources in our low-resource setting.
+In contrast, M2M100 performed well despite these constraints.
+We hypothesize that in higher computational resource settings, LLMs could outperform all other methods, leveraging their capacity for nuanced language understanding when adequately pre-trained and resourced.
+Nonetheless, to validate the reproducibility and robustness of our methods, we conducted additional experiments with Bambara—a language from a different linguistic family.
+The results of this experiment are detailed in Section 7.
+7 Further Experiment with other Languages
+After obtaining promising results for Zarma using LLM and MT-based approaches, we conducted further experiments to validate the reproducibility of these methods—using the M2M100 and MT5 models.
+We selected the Bambara language for this experiment because it belongs to a different linguistic family, which allows us to evaluate the performance of the approaches on a language outside the Nilo-Saharan family.
+We used the Bayelemabaga dataset (Vydrin et al., 2022)—split into 80% 10% 10%—for this experiment.
+The same data preparation process described in the methodology section was followed; however, we excluded any human-annotated data to focus only on spelling correction performance.
+The results are presented in Table 7.
+The Bambara experiment demonstrated that the MT-based approach outperformed the LLMs-based one regarding word-level correction metrics.
+The MT-based approach achieved a detection rate of 94.64% and a suggestion accuracy of 68.18%.
+In contrast, the LLMs-based approach detected 87.45% of errors and suggested corrections with 52.91% accuracy.
+The results from the Bambara experiment highlight the potential of both LLMs and MT models to improve GEC for low-resource languages significantly.
+However, they also emphasize the necessity for continued expanding and diversifying training datasets and pre-training for LLMs.
+8 Conclusion and Future Work
+We presented a comparison of rule-based, LLM-based, and MT-based methods for Zarma GEC.
+Our results provide answers to our research questions.
+Regarding RQ1, we found that MT models (M2M100) outperform both rule-based systems and LLMs like Gemma and MT5 in Zarma GEC.
+For RQ2, we observed that rule-based methods excel at simple spelling correction but fail at context-level errors, whereas MT models handle both effectively.
+Finally, for RQ3, our experiments with Bambara confirm that the MT-based approach is replicable and effective for other low-resource West African languages.
+This research establishes a foundation for future GEC development in other low-resource languages.
+The approaches covered in this study are starting point references and can be further improved using sophisticated methods.
+Future work includes hybrid models that combine rule-based precision with MT and LLM adaptability, data augmentation to improve performance, continuous learning mechanisms, resource optimization, and cross-lingual transfer learning to support a wider range of languages.
+Models Parameters Training Details
+QLoRA Epochs Batch Size GPU Used Lr
+Gemma 2b 2 billion Applied 2 8 P100 $2\times 10^{-4}$
+MT5-small 300 million Not Applied 10 16 T4x2 $2\times 10^{-5}$
+M2M100 418 million Not Applied 2 32 P100 $2\times 10^{-5}$
+Table 8: Training settings for the models, including epochs, batch size, and optimizer details.
+Limitations
+Rule-Based Method
+Rule-based approaches rely on predefined linguistic rules and dictionaries, which limit their adaptability.
+They do not handle new or unseen words well.
+In zero-shot scenarios, where the system encounters words or structures beyond its dictionary, it often fails to correct or detect errors.
+Additionally, the creation and maintenance of linguistic resources is difficult for low-resource languages, which often lack standardized orthography (Scannell, 2007).
+LLMs Method
+LLMs—Gemma 2b and MT5-small—must rely on diverse training data to capture linguistic nuances.
+Yet, existing LLMs are often trained on data from high-resource languages, which reduces performance in languages like Zarma.
+A lack of training samples in these languages causes the model to produce lower-quality suggestions (Bender et al., 2021).
+In addition, fine-tuning LLMs requires significant computational resources, which can be a challenge in low-resource settings.
+MT Method
+MT models—M2M100 in our studies—learn from parallel corpora that map incorrect text to correct text.
+In low-resource languages, such parallel data is often lacking or incomplete (Tiedemann, 2020).
+Even though multilingual models can leverage patterns from other languages, the absence of sufficient in-language data can lead to contextually imprecise corrections.
+Moreover, these models require targeted fine-tuning and periodic updates as language use evolves.
+Synthetic Data Generation
+Our synthetic data generation process—while designed to mimic common human errors based on an initial study with native speakers (as detailed in Section 4)—has some limitations.
+The corruption script may not capture the full nuanced distribution of errors that humans make—particularly complex grammatical or semantic errors that are not easily programmable by rule-based mechanisms.
+This could lead to models being better prepared for certain types of synthetic errors than for the true wide range of errors encountered in real-world Zarma texts.
+Consequently, performance on naturally occurring errors might differ from that observed on our synthetic test sets.
+Acknowledgments
+This research was partially supported by the computer science department of Ashesi University, and the Cite de l'Innovation of Niger.
+Appendix A Errors Being Addressed
+In this section, we explain the types of errors GEC methods address.
+We categorize the errors into two main types: word-level correction (spellchecking) and context-level correction.
+The context-level correction is further divided into logical errors and sentence improvement.
+Below, we define each error type and provide examples in Zarma with English translations to illustrate them.
+A.1 Word-Level Correction
+Word-Level correction involves identifying and correcting typographical errors in individual words.
+These errors are usually due to misspellings, incorrect usage of characters, or typographical mistakes.
+Such errors typically do not alter the grammatical structure of the sentence but affect readability and lexical accuracy.
+  • Example:
+    • Incorrect Zarma: Sintina gaa Irikog na beena da ganda taka.
+    • English Translation (Incorrect): In the beginning God [misspelled form of God] created the heaven and the earth.
+    • Correct Zarma: Sintina gaa Irikoy na beena da ganda taka.
+    • English Translation (Correct): In the beginning God created the heaven and the earth.
+    • Explanation: The word ``Irikog'' is a misspelling of ``Irikoy'' (God). Spellchecking corrects this specific lexical error.
+A.2 Context-Level Correction
+Context-level correction involves errors that go beyond individual words and affect the overall structure, coherence, and meaning of the sentence.
+These errors require an understanding of the surrounding context to identify and rectify.
+We categorize these errors into logical errors and sentence improvement.
+A.2.1 Logical Errors
+Logical errors are inconsistencies in sentence meaning often caused by incorrect verb conjugations, subject-verb agreement issues, or the misuse of grammatical markers such as tense or aspect markers.
+These errors affect the semantic coherence of the sentence and require contextual understanding to correct.
+  • Example:
+    • Incorrect Zarma: Souba, Ay koy Niamey.
+    • English Translation (Incorrect): Tomorrow, I go to Niamey.
+    • Correct Zarma: Souba, Ay ga koy Niamey.
+    • English Translation (Correct): Tomorrow, I will go Niamey.
+    • Explanation: The time indicator ``Souba'' means ``tomorrow,'' which requires a future tense.
+The original sentence incorrectly uses ``koy'' (go/am going), which is present tense.
+The correction introduces the future tense marker ``ga'' before the verb ``koy'' to align the verb tense with the future time reference, resolving the logical inconsistency.
+A.2.2 Sentence Improvement
+Sentence improvement focuses on enhancing the overall readability, precision, naturalness, or stylistic quality of a sentence.
+While the original sentence may already be grammatically correct, improvements aim to make it flow better, be more concise, or adhere more closely to common usage patterns.
+  • Example:
+    • Original Zarma: Albarka ma te ni se teŋ kaŋ ni goy.
+    • English Translation: Blessings be to you for the work that you did.
+    • Improved Zarma: Ni goyom albarka.
+    • English Translation: Blessings on your work / Your work is blessed.
+    • Explanation: The original sentence is grammatically correct but might be perceived as slightly formal or verbose in some contexts.
+The improved version offers a more concise and idiomatic expression common in Zarma for acknowledging good work.
+Appendix B Evaluators Recruitment Process
+To ensure a thorough and culturally appropriate manual evaluation of our GEC methods, we recruited five native Zarma speakers to serve as evaluators, who also assisted in developing the corruption script for synthetic data (Section 4).
+The primary criterion for selection was their fluency in Zarma—their ability to read and write proficiently in the language.
+The evaluators were selected through community outreach in Niamey, Niger.
+All selected evaluators had completed secondary education or higher.
+Prior to the evaluation, we conducted a training workshop to familiarize the evaluators with the objectives of the study, the evaluation protocols, and the scoring guidelines.
+During the evaluation process, each evaluator independently assessed a set of 300 sentences that had been manually corrupted to include logical and grammatical errors.
+They reviewed the corrections made by each GEC method and assigned scores based on accuracy and fluency, where:
+  • 1 indicates a poor correction with significant errors remaining or introduced.
+  • 5 indicates an excellent correction with all errors appropriately addressed and the sentence reading naturally.
+All evaluators gave informed consent to participate in the study and were compensated for their time and contributions.
+Appendix C Potential Applications
 Our team visited Niamey to present our work to the local Zarma community and gather their feedback.
@@
 The GEC tool can support these efforts by providing a means for transcribing and publishing grammatically accurate texts.
-8 Conclusion and Future Work
-We presented a comparison of rule-based, LLM-based, and MT-based methods for Zarma GEC.
-The MT-based approach performed best in both controlled and zero-shot scenarios.
-We confirmed these findings by testing with Bambara language.
-This research establishes a foundation for future GEC development in other low-resource languages.
-Future work includes hybrid models that combine rule-based precision with MT and LLM adaptability, data augmentation to improve performance, continuous learning mechanisms, resource optimization, and cross-lingual transfer learning to support a wider range of languages.
-9 Limitations
-9.1 Rule-Based Method
-Rule-based approaches rely on predefined linguistic rules and dictionaries, which limit their adaptability.
-They do not handle new or unseen words well.
-In zero-shot scenarios, where the system encounters words or structures beyond its dictionary, it often fails to correct or detect errors.
-Additionally, the creation and maintenance of linguistic resources is difficult for low-resource languages, which often lack standardized orthography (Scannell, 2007).
-9.2 LLMs Method
-LLMs—Gemma 2b and MT5-small—must rely on diverse training data to capture linguistic nuances.
-Yet, existing LLMs are often trained on data from high-resource languages, which reduces performance in languages like Zarma.
-A lack of training samples in these languages causes the model to produce lower-quality suggestions (Bender et al., 2021).
-In addition, fine-tuning LLMs requires significant computational resources, which can be a challenge in low-resource settings.
-9.3 MT Method
-MT models—M2M100 in our studies—learn from parallel corpora that map incorrect text to correct text.
-In low-resource languages, such parallel data is often lacking or incomplete (Tiedemann, 2020).
-Even though multilingual models can leverage patterns from other languages, the absence of sufficient in-language data can lead to contextually imprecise corrections.
-Moreover, these models require targeted fine-tuning and periodic updates as language use evolves.
-Appendix A Further Experiment with other Languages
-After obtaining promising results for Zarma using LLM and MT-based approaches, we conducted further experiments to validate the reproducibility of these methods—using the M2M100 and MT5 models.
-We selected the Bambara language for this experiment because it belongs to a different linguistic family, which allows us to evaluate the performance of the approaches on a language outside the Nilo-Saharan family.
-We used the Bayelemabaga dataset (Vydrin et al., 2022) for this experiment.
-The same data preparation process described in the methodology section was followed; however, we excluded any human-annotated data to focus only on spelling correction performance.
-The results are presented in Table 6.
-Methods AE Evaluation
-Common ML Metrics Spelling Correction
-GLEU Detection Correction
-MT5-small 0.5713 0.643 87.45% 52.91%
-M2M100 0.7293 0.851 94.64% 68.18%
-Table 6: Performance Metrics for Bambara
-The Bambara experiment demonstrated that the MT-based approach outperformed the LLMs-based one regarding word-level correction metrics.
-The MT-based approach achieved a detection rate of 94.64% and a suggestion accuracy of 68.18%.
-In contrast, the LLMs-based approach detected 87.45% of errors and suggested corrections with 52.91% accuracy.
-The results from the Bambara experiment highlight the potential of both LLMs and MT models to improve GEC for low-resource languages significantly.
-However, they also emphasize the necessity for continued expanding and diversifying training datasets.
-Appendix B Errors Being Addressed
-In this section, we explain the types of errors GEC methods address.
-We categorize the errors into two main types: word-level correction (spellchecking) and context-level correction.
-The context-level correction is further divided into grammar errors, logical errors, and sentence improvement.
-Below, we define each error type and provide examples to illustrate them.
-B.1 Word-Level Correction (Spellchecking)
-Word-Level correction involves identifying and correcting typographical errors in individual words.
-These errors are usually due to misspellings, incorrect usage of characters, or typographical mistakes.
-  • Example:
-    • Incorrect: Sintina gaa Irikog na beena da ganda taka.
-    • Correct: Sintina gaa Irikoy na beena da ganda taka.
-B.2 Context-Level Correction
-Context-level correction involves errors that go beyond individual words and affect the overall structure and meaning of the sentence.
-We categorize these errors into logical errors and sentence improvement.
-B.2.1 Logical Errors
-Logical errors are inconsistencies in sentence meaning that arise from incorrect verb conjugations, subject-verb agreement, or the misuse of grammatical markers.
-Unlike grammatical errors, logical errors affect the semantic coherence of a sentence.
-  • Example:
-    • Incorrect: Souba, Ay koy Niamey. (The time indicator "Souba" means "tomorrow," but the verb "koy" indicates present tense.)
-    • Correct: Souba, Ay ga koy Niamey. (The future tense marker "ga" matches the time indicator "Souba.")
-B.2.2 Sentence Improvement
-Sentence improvement focuses on enhancing readability, precision, and style. While the sentences may already be grammatically correct, improvements ensure better flow and clarity.
-  • Example:
-    • Original: I girbi honkuna i tun be.
-    • Improved: I ga girbi suba.
-Appendix C Evaluators Recruitment Process
-To ensure a thorough and culturally appropriate manual evaluation of our GEC methods, we recruited five native Zarma speakers to serve as evaluators.
-The primary criteria for selection was their fluency in Zarma—their ability to read and write proficiently in the language.
-The evaluators were selected through community outreach in Niamey, Niger.
-All selected evaluators had completed secondary education or higher.
-Prior to the evaluation, we conducted a training workshop to familiarize the evaluators with the objectives of the study, the evaluation protocols, and the scoring guidelines.
-During the evaluation process, each evaluator independently assessed a set of 300 sentences that had been manually corrupted to include logical and grammatical errors.
-They reviewed the corrections made by each GEC method and assigned scores based on accuracy and fluency, where:
-  • 1 indicates a poor correction with significant errors remaining or introduced.
-  • 5 indicates an excellent correction with all errors appropriately addressed and the sentence reading naturally.
-All evaluators provided informed consent to participate in the study and were compensated for their time and contributions.
-Appendix D Training Details
-Models Parameters Training Details
-QLoRA GPU Used Lr
-Gemma 2b 2 billion Applied P100 $2\times 10^{-4}$
-MT5-small 300 million Not Applied T4x2 $2\times 10^{-5}$
-M2M100 418 million Not Applied P100 $2\times 10^{-5}$
-Table 7: Training Settings for the models
-Appendix E Other Figures
-Figure 2: Images of the different GEC tool interfaces. The rule-based on the left and the other approaches on the right